### v1

# Optimizing Token Choice for Code Watermarking: A RL Approach

Abstract
The need for detecting LLM-generated code necessitates watermarking systems capable of operating within its highly structured and syntactically constrained environment. To address this, we introduce CodeTracer, an innovative adaptive code watermarking framework underpinned by a novel reinforcement learning training paradigm. At its core, CodeTracer features a policy-driven approach that utilizes a parameterized model to intelligently bias token choices during next-token prediction. This strategy ensures that embedded watermarks maintain code functionality while exhibiting subtle yet statistically detectable deviations from typical token distributions. To facilitate policy learning, we devise a comprehensive reward system that seamlessly integrates execution feedback with watermark embedding signals, balancing process-level and outcome-level rewards. Additionally, we employ Gumbel Top-k reparameterization to enable gradient-based optimization of discrete watermarking decisions. Extensive comparative evaluations demonstrate CodeTracer’s significant superiority over state-of-the-art baselines in both watermark detectability and the preservation of generated code’s functionality.
1 Introduction
The unprecedented capabilities of large language models (LLMs) in code generation have introduced critical challenges for intellectual property protection and code attribution [27, 1]. As AI systems produce increasingly sophisticated code that is difficult to distinguish from human-written, the need for reliable code tracing approaches has become urgent [26]. Code watermarking, a technique historically vital for software protection, now serves as the cornerstone for identifying and verifying machine-generated code origins [9, 11, 4]. Unlike natural language, code presents unique watermarking challenges due to its strict structural and semantic requirements: even slight modifications can completely break functionality, a problem less prevalent in text watermarking [13].
LLM generation is an autoregressive process composed of sequential next-token predictions [20]. LLM watermarking biases these predictions toward token choices that subtly differ from normal selections, creating detectable statistical patterns. Many LLM watermarking schemes typically rely on a “green-red list” methodology [9]. This technique biases token generation toward specific green (preferred) tokens and suppressing red (avoided) tokens, creating detectable statistical patterns without significantly degrading output quality. For natural language, these vocabulary partitions can be determined randomly [9, 13] with minimal impact on coherence. Code watermarking faces unique challenges: random token selection often breaks functionality since certain tokens are syntactically essential, while uniform watermarking ignores structural constraints where some positions cannot tolerate modifications.
Therefore, we identify our core challenge: how to intelligently determine optimal watermark insertion points and select semantically reasonable token choices that maintaining statistical detectability patterns while preserve code functionality?
Traditional code watermarking approaches suffer from significant limitations: they operate as post-generation transformations on completed code, making them incompatible with the causal generation process of LLMs. Additionally, these methods require labor-intensive, language-specific transformation rules that must be manually crafted for each programming language [8, 24, 14]. Recent efforts to incorporate watermarks during LLM code generation show promise but remain impractical in real-world scenarios, as they require access to supplementary prompts and model information to compute critical values like entropy during detection [13, 6]. We argue that an effective solution requires a model-based approach that learns the necessary programming knowledge during training, enabling intelligent watermarking decisions that adapt to syntactic and semantic constraints with minimal context.
In this work, we propose CodeTracer, an adaptive code watermarking framework built on a novel reinforcement learning (RL) paradigm. Our approach features a parameterized watermark model that intelligently identifies optimal positions and biases token selection during the code generation process. Reinforcement learning provides an ideal solution for our task as it eliminates the need for pre-existing watermarked code examples. This is crucial because creating such training examples presents a fundamental paradox: we cannot determine whether a code snippet is “watermarked” without first defining the watermark model itself.
Challenges with RL.
Designing a learnable code watermarking RL framework presents significant challenges. Two critical difficulties emerge in this process. First, effective reward mechanisms are fundamental to guiding the RL process, requiring careful integration of available feedback signals. Second, the watermarking procedure necessitates making discrete decisions about code tokens, creating a non-differentiable process that is inherently difficult to optimize. To address these challenges, we implement a dual-component reward system: execution feedback that penalizes functionally incorrect code, complemented by watermark embedding signals consisting of both immediate process rewards for successful watermarked token selection and statistical outcome rewards that employ metrics like z-score to comprehensively measure detectability. Additionally, we overcome the non-differentiability constraint through Gumbel Top-k reparameterization [22], enabling gradient-based optimization by allowing differentiable selection of tokens.
Contributions.
Our work makes several key contributions: (i) We introduce CodeTracer, an adaptive watermarking framework that intelligently embeds watermarks during LLM code generation; (ii) We develop a novel RL pipeline for code watermarking, combines execution feedback with dual watermark signals, and enables differentiable optimization for the whole pipeline; and (iii) Our evaluations across multiple datasets demonstrate CodeTracer’s superior watermark detection capabilities while maintaining code functionality.
2 Related Works
LLM Watermarking.
LLMs’ sophisticated content generation capabilities create challenges for authenticity verification [26, 19, 4]. Recent watermarking approaches embed signatures during the token sampling process by modifying logits or altering the sampling procedure [9, 11]. A notable example is the green-red watermarking scheme [9], which partitions vocabulary tokens into lists and biases selection toward “green” tokens, enabling statistical detection. [23] introduces a RL-based approach for watermark embedding but their co-training scheme may cause unexpected behavior for LLMs. However, these methods struggle with low-entropy conditions common in code generation [13].
Code Watermarking.
Code watermarking imposes unique challenges due to programming languages’ strict syntactic and semantic requirements. Traditional code watermarking alters existing code through format changes or control flow restructuring [8, 17, 14, 24, 16, 4]. Recent LLM-focused approaches use entropy distributions [13, 15] or type predictors [6] to guide watermark insertion. However, these techniques typically require privileged access to LLM parameters, generation probabilities, or original prompts, limiting their practical deployment. CodeTracer instead adaptively embeds watermarks during generation without such prerequisites.
Rule-Based RL and GRPO.
Rule-Based Reinforcement Learning enhances LLMs by integrating explicit rules into the learning process, demonstrating reduced vulnerability to reward hacking [7, 12]. These approaches leverage simple, verifiable rules that ground model outputs, showing particular effectiveness on structured tasks like coding and mathematical reasoning [25]. Deepseek-R1 [7] exemplifies this approach through Group Relative Policy Optimization (GRPO), which significantly reduces computational overhead while improving complex reasoning capabilities. GRPO generates multiple candidate responses for a given prompt, then calculates advantages by normalizing rewards relative to group statistics, eliminating the need for trained critic models in value estimation. Our CodeTracer framework adopts this GRPO scheme with rule-based rewards for computational efficiency.
3 Preliminaries
LLM Watermarking.
In a code generation task, large language model receives a sequence of tokens $\mathbf{x}=[x_{1},x_{2},\ldots,x_{n}]$ as the input prompt, and generates output $\mathbf{y}=[y_{1},y_{2},\ldots,y_{m}]$ including the code snippet. At every generation step $t$, the LLM $\pi_{\theta}$ receives the input prompt $\mathbf{x}$ and previous generated token to compute the logit vector $\mathbf{l}=[l_{1},l_{2},\ldots,l_{|\mathcal{V}|}]$ for token $t$, cross all tokens in the vocabulary $\mathcal{V}$, where each $l_{j}$ corresponds to token $v_{j}\in\mathcal{V}$.. The logit vector is then applied with a softmax opreator to transform it into a probability distribution where token at position $y_{t}$ is sampled from. To detect AI generated content, in-generation watermarking [9] adjust the original LLM generation policy by adding constant biases on a sublist of vocabulary that are selected by a pseudorandom function, resulting in a watermarked code sequence $\mathbf{\tilde{y}}=[\tilde{y}_{1},\tilde{y}_{2},\ldots,\tilde{y}_{m^{\prime}}]$. The statistical difference in the token distribution within/out the sublists are then utlized to detect the AI-generated output.
Challenge of LLM Code Watermarking.
However, LLM watermark faces serve performance degration in the code because of unique challenges distinct from natural text watermarking. Unlike natural language, code is highly structured with a delicate syntax where small modifications can drastically alter functionality or render the code inoperable. This inherent rigidity creates significant constraints for watermarking techniques. First, watermark position selection is particularly critical in code, as certain positions are immutable. Unlike natural text where minor alterations rarely affect meaning, code contains strict structural elements that cannot be modified. For example, altering the def keyword in a Python function definition (e.g., changing def process_data(data): to func process_data(data):) inherently breaks syntax. Any watermarking approach must identify and avoid these critical positions to maintain functional code. Second, watermark token choice must respect contextual constraints. Even at positions where modification is possible, the replacement tokens must be semantically and syntactically appropriate. For instance, if a model generates an assignment like status = "active" but a watermarking scheme forces a substitution of "active" with an incompatible token such as the keyword class, the resultant status = class would cause a syntax error. This contextual sensitivity significantly limits the vocabulary available for watermarking while preserving code functionality.
Why RL for $\pi_{\phi}(a|\mathbf{c})$ training?
Reinforcement learning provides an ideal framework for this task for three key reasons: (i) Learning complex rules implicitly: The diversity of programming languages makes manually defining watermarking rules impractical. RL can learn syntactic constraints across various code structures through interaction rather than exhaustive specification. (ii) Overcoming the absence of supervised data: SFT is impossible due to a circular dependency—we cannot create "gold standard" watermarked examples without first defining the watermark model itself. RL enables learning optimal strategies without requiring predefined examples. (iii) Leveraging clear feedback signals: Both watermark effectiveness and code functionality provide quantifiable rewards. Our system utilizes rule-based feedback mechanisms that have proven effective for code tasks [7, 12], making RL well-suited for optimizing within code generation constraints.
4 Learn to Watermark
We formulate the LLM code watermarking problem within a reinforcement learning (RL) framework. This section first introduces the watermark policy, then outlines the process of training this policy through RL, and finally addresses specific design challenges along with our proposed solutions.
4.1 Watermarked Policy
We focus on a single step in the token generation process, specifically one next token prediction. Watermark model $\pi_{\phi}(a|\mathbf{c})$ operates based on the current context $\mathbf{c}$ (comprising part of input $\mathbf{x}$ and previously generated tokens in $\mathbf{y}$ up to the current position with a fixed number of tokens). This model outputs an action $a=(w,G)$ consisting of two components: a binary decision $w\in\{0,1\}$ indicating whether to apply watermarking at this position, and a green token list $G\subset\mathcal{V}$ for preferred tokens.
The watermarked token generation process combines the original LLM $\pi_{\theta}$ with the watermark model $\pi_{\phi}$ and get watermarked policy $\pi_{\theta\oplus\phi}$, computing the modified logits as:
$\tilde{l}_{j}=l_{j}+w\cdot\delta\cdot\mathbbm{1}_{v_{j}\in G}$ (1)
where $\delta$ represents a bias added exclusively to green tokens when watermarking is active ($w=1$), and $\mathbbm{1}_{v_{j}\in G}$ is an indicator function equaling 1 if token $v_{j}\in G$ and 0 otherwise. The watermark strength is controlled by two constraints: the size of the green token set where $|G|=\gamma|\mathcal{V}|$ for a predefined parameter $\gamma\in(0,1)$, and the bias value $\delta$. The remaining tokens form a “red” group $R=\mathcal{V}\setminus G$. This policy-driven, selective approach dynamically controls both watermark placement and vocabulary partitioning used for biasing.
With this formulation, the policy strategically biases token choices through the green set $G$ and controls watermark application via the binary decision $w$. The core challenge lies in training the parameters $\phi$ of the watermark policy model. In the following subsection, we present the complete watermark pipeline for both watermark generation and detection using the combined policy $\pi_{\theta\oplus\phi}$, followed by our training methodology that optimizes the watermark policy $\pi_{\phi}$ in next section.
Generation
During code generation, our framework CodeTracer applies the previously defined framework sequentially. The watermarked policy $\pi_{\theta\oplus\phi}(\mathbf{\tilde{y}}|\mathbf{x})$ takes exactly same input with LLM $\pi_{\theta}$ but generate the biased output sequences $\mathbf{\tilde{y}}=[\tilde{y}_{1},\tilde{y}_{2},\ldots]$ based on biased logits $\tilde{l}_{j}$ at each step.
Detection.
For detection, CodeTracer analyzes a potentially corrupted sequence $\mathbf{s}=[s_{1},s_{2},\ldots,s_{T^{\prime}}]$ of length $T^{\prime}$. The watermark policy $\pi_{\phi}$ reconstructs its decisions independently, without requiring the LLM $\pi_{\theta}$. This reconstruction process determines two critical pieces of information for each token: first, whether watermarking was applied at that position ($w=1$ or $w=0$), and second, which tokens belonged to the green list $G$ if watermarking was applied. By applying this reconstruction, we identify the subset of tokens $s_{w=1}$ where watermarking was active, yielding a total count of watermarked positions $T=|\{s_{w=1}\}|$. Among these watermarked positions, we then count how many tokens fall into their respective green sets, denoted as $N_{G}=|\{s_{w=1,s\in G}\}|$. A one-proportion $z$-test assesses statistical significance:
$z=\frac{N_{G}-T\gamma}{\sqrt{T\gamma(1-\gamma)}}.$ (2)
A significantly large positive $z$-score indicates watermark presence, demonstrating that watermarked tokens appear in their respective green sets more frequently than would occur by random chance.
Figure 1: CodeTracer: A framework for LLM code watermarking through selective token biasing. The diagram illustrates the end-to-end pipeline where a trainable watermark model works alongside an LLM to embed statistical patterns in code, with a reward system that optimizes for both code functionality and watermark detectability.
4.2 How to Learn the Watermarked Policy?
As shown in Figure 1, to train the watermark model $\pi_{\phi}(a|\mathbf{c})$ in the watermarked policy $\pi_{\theta\oplus\phi}$, we employ GRPO as our RL framework. We selected GRPO for its computational efficiency and demonstrated effectiveness when integrating rule-based rewards on coding and mathematical tasks [7]. In a standard GRPO training scheme, three components are essential: a training policy, a reference policy, and reward functions. Our training pipeline uses $\pi_{\theta\oplus\phi}$ as the training policy rather than $\pi_{\phi}$ alone, since the complete generation process requires the coordinated operation of both the LLM and the watermark model. Importantly, we freeze the LLM parameters $\theta$ during training and optimize only watermark model parameters $\phi$, ensuring a portable and efficient watermark solution.
4.3 Policy Design
The implementation of our watermarked policy $\pi_{\theta\oplus\phi}$ presents fundamental technical challenges due to the non-differentiable nature of several components in Equation (1). Specifically, we implement the watermark model $\pi_{\phi}(a|\mathbf{c})$ as a transformer-based architecture. While the transformer outputs are continuous, we need to convert them to discrete outputs while maintaining differentiability. The watermark binary decision output $w$ and green token list assignment $G$ involve discrete selection processes that create optimization barriers in the neural network training pipeline.
Straight-Through Estimation.
To enable gradient flow through the discrete decision processes for $w$, we apply straight-through estimation techniques. The watermark model transformer architecture first outputs a $|l_{\phi}|+1$ dimensional vector, which we represent as $(w_{\phi},l_{\phi})$. To derive our desired output $a=(w,G)$ from $(w_{\phi},l_{\phi})$, we implement the straight-through estimations:
$w=\mathbbm{1}_{w_{\phi}>0.5}+\sigma(w_{\phi})-\text{sg}(\sigma(w_{\phi})).$ (3)
During the forward pass, this formulation behave like the discrete operations $\mathbbm{1}[w_{\phi}>0]$. However, during backpropagation, gradients flow through the continuous relaxations $\sigma(w_{\phi})$, with $\text{sg}(\cdot)$ ensuring that only the gradient of the relaxation (and not their forward values) affects the computation. This technique enables end-to-end training despite the discrete nature of watermarking decisions.
Gumbel-Top-k.
To overcome the challenges of discrete token selection for the green list $G$, we develop a Gumbel-Top-k sampling approach for our token assignment process. This technique provides a differentiable pathway from $l_{\phi}$ to $G$ while preserving the discrete characteristics necessary for effective watermarking. For a given parameter $\gamma\in(0,1)$, we select $k=\lfloor\gamma|\mathcal{V}|\rfloor$ tokens for the green list. Given logits $\mathbf{l}_{\phi}$ from the watermark model $\pi_{\phi}$, we apply Gumbel noise to create a continuous relaxation of the discrete selection:
$\mathbf{g}=\mathbf{l}_{\phi}+(-\log(-\log(\mathbf{u})))$ (4)
where $\mathbf{u}\sim\text{Uniform}(0,1)^{|\mathcal{V}|}$ represents uniform random noise across the vocabulary space. The green list $G$ is then determined by selecting the top-$k$ tokens according to these perturbed logits:
$G=\text{arg top-{k}}(\mathbf{g})$ (5)
To enable gradient flow through this discrete selection process, we implement a straight-through estimator for the indicator variable $\mathbf{l}_{G}\in\{0,1\}^{|\mathcal{V}|}$ that represents membership in the green list:
$\mathbf{l}_{G}=\mathbbm{1}_{v\in G}+\mathcal{S}(\mathbf{g})_{v}-\text{sg}(\mathcal{S}(\mathbf{g})_{v})$ (6)
where $\mathcal{S}(\mathbf{g})_{v}$ represents the Gumbel-Softmax-based continuous relaxation of the top-$k$ selection for token $v$, and $\text{sg}(\cdot)$ is the stop-gradient operator. During forward passes, we use hard discrete token selection based on token rankings from the Gumbel-perturbed logits, while during backward passes, the gradients flow through the softmax-based continuous relaxation. This dual-path approach allows the model to learn effective watermarking strategies while maintaining the discrete token selection.
For the reference policy component in our GRPO framework, we utilize the same architecture as the training policy $\pi_{\theta\oplus\phi}$, establishing a self-referential optimization approach. This design choice enables the watermarking parameters to be refined against their own previous distributions. The approach facilitates gradual refinement of watermark policy while ensuring that successive iterations remain coherent with previous behaviors, ultimately leading to more stable convergence and better performance.
4.4 Reward Function Design
The goal of the reward system is to leverage watermark detection feedback and code execution signals to guide watermarked policy $\pi_{\theta\oplus\phi}$ in selecting suitable watermarking positions ($w$) and making effective token choices for the green list ($G$). To achieve this, we have designed a reward system that comprises a set of reward functions balancing the objectives crucial for code watermarking.
Outcome-based Execution Reward $R_{1}$.
Our execution reward $R_{1}$ implements a binary evaluation of the watermarked code’s functional integrity through test case performance. We define:
[…]
This straightforward formulation provides clear feedback on code correctness: successful execution of all test cases yields a positive reward, while any compilation error or runtime failure incurs a substantial penalty. By incorporating this binary signal into our reward function, we establish a strict constraint that compels the watermark policy to preserve the code’s original functionality.
Outcome-based Watermark Reward $R_{2}$.
The watermark detection reward $R_{2}$ quantifies the statistical detectability of embedded watermarks within generated code sequences using a saturated function based on the statistical z-score. For each complete sequence $s$, we implement:
$R_{2}(s)=\begin{cases}1,&\text{if }z(s)\geq 3\\ \frac{z(s)}{3},&\text{if }0<z(s)<3\\ 0,&\text{if }z(s)\leq 0\end{cases}$ (8)
where $z(s)$ represents the statistical z-score from our detection algorithm. This saturated reward formulation establishes clear boundaries: a z-score of 3 or higher (indicating strong statistical significance) receives the maximum reward of 1, while non-positive z-scores receive no reward. Between these thresholds, the reward scales linearly with the z-score. This design incentivizes the watermark policy to achieve statistically significant detection levels without wasting computational resources on unnecessarily strong signals.
Process-based Watermark Reward $R_{3}$.
In our implementation, we find that outcome-based rewards alone are insufficient for effective policy learning, as they assign equal advantage to both red and green tokens when the overall sequence succeeds. To provide more granular guidance, we introduce a process-based reward $R_{3}$ that evaluates each token individually, offering immediate feedback on token selection decisions. This approach is straightforward to implement as we can directly check whether each generated token $s_{t}$ is green, red, or non-watermarked based on the policy’s action $a_{t}=(w_{t},G_{t})$ at the $t$-th token:
[…]
This simple yet effective reward structure explicitly encourages the selection of green tokens during watermarking while penalizing red token selection. Non-watermarked positions receive a neutral reward, focusing the learning signal on positions where watermarking is applied. By providing direct feedback at each token generation step, the process-based reward creates a much clearer learning landscape compared to sequence-level outcome rewards alone. This token-by-token guidance significantly accelerates policy convergence and leads to more effective watermarking strategies.
Table 1: Performance (%) comparison of text detection methods. Methods marked with work in an impractical scenario with original LLM and prompts. Bold values indicate best performance within each category.
Category Dataset HumanEval MBPP
Method / Metric Pass@1 Pass@10 AUROC TPR Pass@1 Pass@10 AUROC TPR
No Watermark Base 65.42 79.17 - - 43.35 51.65 - -
Post-hoc Methods logp(x) 65.42 79.17 47.59 4.27 43.35 51.65 47.77 6.40
LogRank 65.42 79.17 47.66 1.82 43.35 51.65 48.76 7.80
DetectGPT 65.42 79.17 51.12 9.15 43.35 51.65 46.15 3.60
GPTZero 65.42 79.17 52.00 5.50 43.35 51.65 41.10 2.80
Oracle Watermarks SWEET 61.77 75.45 83.10 42.68 41.01 48.35 84.40 32.00
Practical Watermarks WLLM 58.05 70.35 70.17 20.73 39.66 47.22 76.44 27.80
EXP-edit 59.29 72.41 66.50 25.61 40.16 50.25 51.22 10.60
CodeTracer 60.82 79.72 82.95 46.34 41.45 51.15 79.69 22.60
Integration into Advantage Computation.
Our approach calculates advantages at two complementary levels to optimize the watermarked policy. First, we compute outcome advantages $A_{1}$ from the execution and watermark detection rewards ($R_{1}$ and $R_{2}$) using GRPO’s outcome reward methodology. Second, we derive process advantages $A_{2}$ from the token-level watermark reward ($R_{3}$), normalizing these values across all tokens in all rollouts for each prompt. Our framework categorizes tokens into four distinct types: green tokens (watermarked and included in set $G$), red tokens (watermarked but falling in set $R$), non-watermarked tokens (where decision $w=0$), and non-code text tokens. To maintain focus specifically on code watermarking, we exclude non-code tokens entirely from process advantage calculations.
The final advantage function for each token $s_{t}$ integrates these signals directly:
$A_{\text{total}}(s_{t},a_{t})=A_{1}+A_{2}(s_{t},a_{t}),$ (10)
where $A_{1}$ represents outcome-based advantage and $A_{2}$ process-based advantage for token $s_{t}$ with action $a_{t}$. We then apply a binary mask to eliminate advantage values for any non-code tokens:
$\hat{A}(s_{t},a_{t})=A_{\text{total}}(s_{t},a_{t})\cdot\mathbbm{1}_{\text{is\_code}}(s_{t}).$ (11)
This integrated advantage function balances global sequence-level feedback with targeted token-specific guidance. While the outcome component $A_{1}$ evaluates tokens within the context of overall watermark effectiveness and functional code execution, the process component $A_{2}$ provides immediate direction for token selection decisions.
Ideally, our integrated reward system assigns high advantage to tokens that simultaneously maintain code functionality, contribute to effective watermark detection, and are selected from the green set. This comprehensive approach guides our watermarked policy $\pi_{\theta\oplus\phi}$ to identify optimal watermarking positions and select tokens that preserve functionality while introducing subtle, statistically detectable deviations from normal token distributions. By selectively masking non-code tokens, we concentrate the learning signal exclusively on elements most relevant to code watermarking, substantially improving policy optimization efficiency.
KL Regularization.
To ensure that the watermarked policy $\pi_{\theta\oplus\phi}$ does not deviate excessively from natural code generation patterns, thereby maintaining code stealth and linguistic quality, we incorporate a Kullback-Leibler (KL) divergence penalty into the overall GRPO loss function. This term, $\beta\cdot D_{\text{KL}}(\pi_{\theta\oplus\phi}\parallel\pi_{\text{ref}})$, regularizes the watermarked policy against the reference policy $\pi_{\text{ref}}$.
5 Experiments
Evaluation Dimensions and Metrics.
We evaluate CodeTracer across three critical dimensions: (i) code functionality, (ii) watermark detectability, and (iii) robustness against attacks. For code functionality, we employ the Pass@k metric [3]. Watermark detection performance is measured by AUROC and TPR@5%FPR metrics.
Comparative Baselines.
For a comprehensive comparison, we compare CodeTracer against established post-hoc detection methods including logp(x), LogRank [5], DetectGPT [18], and GPTZero [21]. Moreover, we compare with active watermarking approaches such as WLLM [9] and EXP-edit [11]. We also include SWEET [13] serves as a reference baseline. However, to be noted, SWEET needs to redo the forward during the detection phase, which adopts a different setting from other baselines. A detailed discussion can be found in Appendix C.2.
Technical Implementation.
We primarily utilize OpenCoder-1.5B-Instruct [Huang2024OpenCoderTO] as the backbone LLM, with additional validation on the larger OpenCoder-8B-Instruct model. Performance is assessed on HumanEval [3] and MBPP [2] for Python code generation. Detailed implementation specifications, including dataset details, baseline configurations, model architectures, and training details are provided in Appendix C. In accordance with established reinforcement learning practices, we initially apply supervised fine-tuning to $\pi_{\phi}$ using proxy signals to establish fundamental code distribution knowledge through a combination of next token prediction and entropy distribution with binary cross entropy loss. Details and other techniques involved can be found in Appendix B.
5.1 Results on Watermark Detection Performance and Code Functionality
We present experiments evaluating CodeTracer’s performance on code functionality and watermark detection with HumanEval and MBPP in Table 1. CodeTracer outperforms other practical watermarking methods across most metrics. It shows minimal code quality degradation (Pass@1 scores of $60.82\%$ on HumanEval and $41.45\%$ on MBPP), while enabling effective detection (AUROC values of $82.95\%$ and $79.69\%$). With the highest TPR ($46.34\%$) on HumanEval among practical watermarks and detection performance comparable to oracle methods, CodeTracer emerges as the leading practical solution on these two datasets.
5.2 Robustness Analysis
Table 2: Robustness evaluation (%) under code modification attacks. Bold values indicate best performance.
Attack Metric WLLM EXP-edit CodeTracer
Original AUROC 70.17 66.50 82.95
TPR 20.73 25.61 46.34
DIPPER AUROC 55.92 51.21 58.42
TPR 12.81 8.54 14.31
Rename AUROC 70.91 62.02 73.36
TPR 20.12 9.76 29.11
Figure 2: Performance for OpenCoder-8B-Instruct.
Figure 3: The training dynamics for CodeTracer-0.
As shown in Table 2, we evaluate CodeTracer’s robustness against DIPPER [10] paraphrasing and variable renaming attacks [13]. While all methods experience performance degradation under attacks, CodeTracer consistently outperforms competing approaches. Against DIPPER attacks, CodeTracer maintains the highest detection capability (AUROC 58.42%, TPR 14.31%). Similarly, under variable renaming attacks, CodeTracer demonstrates superior resilience (AUROC 73.36%, TPR 29.11%), indicating its effectiveness in real-world scenarios with code modifications.
5.3 Further Analysis
Scalability to Larger Models.
To verify the scalability or our approach, we evaluate CodeTracer on the larger OpenCoder-8B-Instruct model. Results are shown in Figure 3. CodeTracer achieves minimal code functionality degradation (71.77% Pass@1 vs. 72.04% for Base) while significantly outperforming other practical watermarking methods in detection capability (78.69% AUROC compared to 65.90% for WLLM and 54.21% for EXP-edit). These results demonstrate CodeTracer’s effectiveness and generalizability across different parameter scales.
Figure 4: Performance (%) comparison with variant
Dataset Performance Metrics
Method Pass@1 AUROC TPR
CodeTracer 60.82 82.95 46.34
CodeTracer-0 58.65 77.50 27.00
Training from Scratch.
To evaluate the effectiveness of pure reinforcement learning without prior knowledge, we introduce CodeTracer-0, a variant trained without SFT initialization. As shown in Figure 3, despite beginning with low detection rewards (approximately 0.3), the model consistently improves throughout RL training, eventually stabilizing around 0.8 for outcome-based watermark detection reward $R_{2}$, which is a proxy for z-score. Table 4 quantifies the final performance: CodeTracer-0 achieves reasonable results with Pass@1 of 58.65%, AUROC of 77.50%, and TPR of 27.00%, though not matching our fully-initialized CodeTracer. These results demonstrate that while good initialization enhances RL performance, our watermarking method remains effective even when trained entirely from scratch.
Figure 5: Computational Efficiency Comparison
Component Resource Metrics
Method Parameters Memory Inference
[…]
Baseline 1.5B Model 1,500M $\sim$3-6GB $\sim$500-800ms
Relative Overhead $10\% <$$15\% <$$0.02\%$
Computation Efficiency
[…]
6 Conclusion
In this paper, we presented CodeTracer, a novel adaptive code watermarking framework that addresses the challenges of watermarking LLM-generated code while preserving functionality. Our policy-driven approach integrates a watermark model with the base LLM to form a composite policy that dynamically decides when to watermark and which tokens to select. By leveraging reinforcement learning through GRPO, our approach intelligently navigates programming language constraints without requiring pre-existing watermarked examples. Our multi-component reward system balances statistical detectability with code integrity through execution feedback and token-level guidance. Evaluations demonstrate that CodeTracer significantly outperforms existing approaches in both detection accuracy and functionality preservation, offering a promising solution for identifying AI-generated code while maintaining its utility.
7 Limitations
While CodeTracer demonstrates significant advantages, certain limitations merit consideration. The reinforcement learning process for training the watermarking policy $\pi_{\theta}$, though optimized with rule-based rewards, can be computationally intensive, especially when applied to extensive codebases or a wide array of programming languages. The evolving nature of adversarial attacks also means that while CodeTracer enhances robustness, ongoing research will be crucial to maintain resilience against future sophisticated attempts at watermark evasion.
The dynamic, per-token decision-making by the watermarked policy $\pi_{\theta\oplus\phi}$, including its determination of green/red list partitions, could also introduce a slight inference latency during code generation compared to simpler, non-adaptive watermarking methods; the impact of this would depend on the specific complexity of the learned policy. Finally, while our use of Gumbel Soft Top-k effectively enables differentiable learning for discrete vocabulary partitioning, the characteristics of this approximation and its potential interaction with highly diverse or unusual coding constructs represent an area for continued detailed analysis.
Appendix A Watermark Generation and Detection Algorithm of CodeTracer
A.1 Watermark Generation
Algorithm 1 presents our watermark generation process. The algorithm embeds watermarks into code sequences while preserving their functionality. It takes as input a prompt $\mathbf{x}$, context window size $c$, and watermarking hyperparameters including green list ratio $\gamma$ and bias $\delta$ that control the strength and detectability of the embedded watermark.
At each timestep $t$, the watermark model processes the context window to determine whether to apply watermarking ($w$) and which tokens to include in the green list ($G$). The algorithm modifies logits selectively based on these decisions, applying bias $\delta$ to green list tokens when $w=1$.
Algorithm 1 CodeTracer Watermark Generation
0: Prompt $\mathbf{x}$, Context window size $c$, Green list ratio $\gamma$, Bias strength $\delta$
0: Watermarked code sequence $\mathbf{\tilde{s}}$
1: Initialize empty sequence $\mathbf{\tilde{s}}$=[]
2:for each generation step $t$ do
3:  Compute base LLM logits $\mathbf{l}=\pi_{\theta}(\mathbf{ctx})$
4:  Compute context window $\mathbf{ctx}=\text{concat}(\mathbf{x},\mathbf{\tilde{s}})[-(c):]$
5:  Compute watermark model outputs $(w_{\phi},\mathbf{l}_{\phi})=\pi_{\phi}(\mathbf{ctx})$
6:  Compute watermark decision $w=\mathbbm{1}_{w_{\phi}>0.5}$
7:  if $w=1$ then
8:   Select green list $G=\text{arg top-}k(\mathbf{l}_{\phi})$ and red list $R=\mathcal{V}\setminus G$, where $k=\lfloor\gamma|\mathcal{V}|\rfloor$
9:   Compute modified logits: $\tilde{l}_{j}=l_{j}+\delta\cdot\mathbbm{1}_{v_{j}\in G}$ for each token $v_{j}$ in vocabulary $\mathcal{V}$
10:  else
11:   Set $\mathbf{\tilde{l}}=\mathbf{l}$ (no modification)
12:  end if
[…]
14:  Append $\tilde{s}_{t}$ to $\mathbf{\tilde{s}}$
15:end for
16:return  $\mathbf{\tilde{s}}$
A.2 Watermark Detection
Algorithm 2 details our statistical watermark detection procedure. Following the framework of [9], we employ hypothesis testing with a key modification, that testing is performed selectively only at positions likely to contain watermarks.
Given a code sequence $\mathbf{s}$, the detector first reconstructs the vocabulary partitions and switch probabilities using the same watermark model from generation. The detection employs a one-sided hypothesis test using the z-statistic to determine whether green tokens appear more frequently than random chance would predict, with a positive z-score above the threshold indicating watermark presence.
Algorithm 2 CodeTracer Watermark Detection
0: Code sequence $\mathbf{s}$, Context window size $c$, Green list ratio $\gamma$, Detection threshold $\tau$
0: Detection result (watermarked/not watermarked/insufficient data)
1: Initialize counters: $\text{watermarked\_positions}=0$, $\text{green\_tokens}=0$
2: Set $\text{positions\_checked}=\{\}$
3:for each position $t$ in $\mathbf{s}$ do
4:  Compute context window $\mathbf{ctx}=\mathbf{s}[\max(0,t-c):t]$
5:  Compute watermark model outputs $(w_{\phi},\mathbf{l}_{\phi})=\pi_{\phi}(\mathbf{ctx})$
6:  Compute watermark decision $w=\mathbbm{1}_{w_{\phi}>0}$
7:  if $w=1$ then
8:   Reconstruct green list $G=\text{arg top-}k(\mathbf{l}_{\phi})$, where $k=\lfloor\gamma|\mathcal{V}|\rfloor$
9:   $\text{watermarked\_positions}++$
10:   if $s_{t}\in G$ then
11:    $\text{green\_tokens}++$
12:   end if
13:   Add $t$ to positions_checked
14:  end if
15:end for
16: Compute z-score: $z=\frac{\text{green\_tokens}-\text{watermarked\_positions}\cdot\gamma}{\sqrt{\text{watermarked\_positions}\cdot\gamma\cdot(1-\gamma)}}$
17:if $z>\tau$ then
18:  return  “watermarked”
19:else
20:  return  “not watermarked”
21:end if
Appendix B Advanced Training And Generation Techniques
B.1 SFT Initialization
While our reinforcement learning (RL) algorithm can train from random initialization, we find that supervised fine-tuning (SFT) provides a significantly more effective starting point. Since watermarked code is unavailable for direct supervision, we instead use entropy and next token prediction of clean code as a proxy. During SFT, the watermark model $\pi_{\phi}$ is trained on two complementary objectives: predicting watermark decisions $w_{\phi}$ to learn entropy distribution patterns across code structures, and optimizing logits $l_{\phi}$ to capture the underlying distribution of tokens in code sequences. This dual-objective approach establishes a strong foundation that encodes both code syntax patterns and appropriate watermarking signals before RL optimization begins.
B.2 Improved Sampling Strategy
We implement a constrained sampling approach during both generation and detection phases to enhance watermark robustness without additional training. Rather than naively selecting the top-$k$ tokens by logit values, we perform a deterministic sorting of logits $l_{\phi}$ and then apply pairwise or tuple-based sampling constraints. This technique preserves the overall token distribution characteristics for unwatermarked positions while creating a more distinct statistical separation between "green" and "red" tokens for watermarked positions. The approach effectively maintains natural code generation quality while improving watermark detection reliability.
B.3 Entropy Regularization
During policy optimization, we observed the watermark model occasionally converging toward degenerate solutions where watermark signals $w$ approach binary extremes (consistently 0 or 1). To mitigate this issue, we incorporate an entropy regularization term in the policy optimization objective:
$L_{\text{entropy}}=-H(\pi_{\phi}(\sigma(w_{\phi})|\mathbf{c})),$ (12)
where $H(\pi_{\phi})$ represents the entropy of the watermark decision distribution. This encourages the model to maintain more balanced probability distributions for watermark decisions, preventing overfitting and improving robustness across diverse code contexts. Additionally, we implement gradient clipping to further stabilize training dynamics.
B.4 Reward Design for Trivial Solutions
Our initial reward design occasionally led to trivial solutions where the model would either apply watermarking too aggressively or too conservatively. To address this, we refined the process-based watermark reward $R_{3}$ with an asymmetric penalty structure:
[…]
where $\alpha>1$ (typically 2-5) creates a stronger penalty for red token selection when watermarking is active. This asymmetric design incentivizes the model to be more selective about when to apply watermarking ($w=1$), only doing so when it has high confidence in successfully biasing toward green tokens. For less certain positions, the model learns to disable watermarking ($w=0$), which receives a neutral reward. This approach significantly reduces false positive detection rates while maintaining high true positive rates.
Parameter Value Description
Watermark Configuration
Watermark bias ($\delta$) 2.0 Logit adjustment magnitude for green tokens
Green list ratio ($\gamma$) 0.5 Proportion of vocabulary in green list
Context window size ($c$) 2 Number of tokens used for watermark decisions
Z-score threshold 4.0 Statistical threshold for watermark detection
Switch threshold 0.5 Threshold for binary watermark decision
Model Architecture
Base model OpenCoder-1.5B-Instruct Foundation model for code generation
Model dimension ($d_{\text{model}}$) 512 Hidden dimension for watermark policy network
Transformer layers 6 Number of transformer encoder layers
Attention heads 8 Multi-head attention heads per layer
Feed-forward dimension 2048 Intermediate dimension in feed-forward networks
Training Configuration
Learning rate 1.0e-5 Initial learning rate for optimization
LR scheduler Cosine with min rate Decay schedule with minimum rate of 0.1
Batch size 8 Sequences per device for training
Gradient accumulation 4 Steps between gradient updates
Training steps 500 Total optimization steps
Warmup ratio 0.03 Proportion of steps for learning rate warmup
KL coefficient ($\beta_{\text{KL}}$) 0.0 Weight for KL divergence regularization
Number of generations 8 Code completions per prompt during training
Temperature 1.0 Sampling temperature for generation
Table 3: Comprehensive hyperparameter settings for the CodeTracer watermarking framework. The parameters are organized by their functional categories: watermark configuration controlling detectability, model architecture defining the watermark policy network structure, training configuration governing the optimization process, and reward configuration balancing the various objectives.
B.5 Limitations of Using LLMs as Watermark Policy Networks
While using pre-trained LLMs directly as watermark models might seem appealing, we identified several critical limitations that justify our specialized architecture. First, distribution alignment presents a fundamental challenge: LLM token distributions naturally align with the target model distribution, causing “green” token lists to predominantly contain frequently used tokens. This similarity makes distinguishing between unwatermarked and watermarked sequences difficult, as both produce similarly high z-scores during detection. Second, computational efficiency concerns arise as large-scale LLMs with billions of parameters introduce significant processing latency compared to our specialized watermark policy, which operates with orders of magnitude fewer parameters while maintaining task-appropriate performance.
Third, context length mismatch creates architectural incompatibilities, as LLMs are optimized for extended contexts (thousands of tokens), while our watermark policy intentionally operates with minimal contextual information (typically just a few tokens) to enable efficient decision-making at each generation step. Fourth, entropy prediction calibration poses practical challenges because LLMs produce poorly calibrated uncertainty estimates for short contexts, particularly at document beginnings, which compromises the reliability of watermark decisions in these critical positions. Our purpose-built watermark policy network addresses these limitations with a lightweight, specialized architecture that achieves superior watermarking performance while maintaining the natural fluency of generated code.
Appendix C Implementation Details
Table 3 presents the comprehensive hyperparameter configuration for our CodeTracer framework, categorized by functional components to facilitate reproducibility and highlight critical design decisions across our experimental setup. For paper writing, we use AI tools to refine the overall expression.
C.1 Datasets
We evaluate on standard code generation benchmarks that span multiple programming languages and tasks. The primary evaluation uses HumanEval [3] and MBPP [2], which provide comprehensive Python programming challenges with associated test cases and reference implementations.
C.2 Baselines
We compare against two categories of methods: post-hoc detection and active watermarking. Post-hoc detection methods preserve original generation and include zero-shot approaches: logp(x), LogRank [5], DetectGPT [18], and GPTZero [21]. Active watermarking methods include WLLM [9] and EXP-edit [11], which operate under the practical constraints. We also include SWEET [13] as a reference, though comparisons are not fair since it requires access to the original LLM and prompts for entropy calculation. We thoroughly tune the hyperparameters for each baseline following their setup.
DetectGPT.
For the DetectGPT implementation, we used T5-3B as our model. Following the original DetectGPT paper and SWEET [18, 13] , we set the span length to 2 words and applied masking to 20% of the text. For each test, we generated 100 perturbations to ensure robust detection.
SWEET.
For the SWEET baseline implementation, we conducted a systematic hyperparameter search for the entropy threshold, exploring values from 0.3 to 1.2 with increments of 0.3. Our experiments revealed optimal performance with an entropy threshold of 1.2 for the HumanEval dataset and 0.3 for MBPP. In detection staga, SWEET needs to use original LLM, prompt and complete generation sequence to compute the entropy. This requirement severely restricts their practical applicability in real-world scenarios. For practical deployment, an ideal solution would depend solely on the code snippet itself, as access to the generation context is typically unavailable. Thus, we only use SWEET as a reference model in our performance comparision.
EXP-edit.
We evaluate EXP-edit [11] following the methodology of [13]. In our experiments, we set temperature=$0.2$ and top-p=$0.95$. We systematically explore block sizes ($20$ tokens), key sequence lengths ($100$), resample sizes ($50$ runs), and edit distance thresholds ($\gamma=0.0$). Through extensive parameter tuning, we determine the optimal configuration: key length $100$, block size $20$, $50$ sampling runs, and detection threshold $0.1$, which balances watermark detection reliability with code functionality.
C.3 Model Architecture Specifications
Our model implements a transformer-based architecture optimized for code analysis and generation. The model uses pre-norm design with layer normalization before the attention and feed-forward computations to enhance training stability. It consists of four main components:
  • Embedding Layer: Combines token embeddings ($W_{\text{te}}\in\mathbb{R}^{|\mathcal{V}|\times d_{\text{model}}}$) and position embeddings ($W_{\text{pe}}\in\mathbb{R}^{c\times d_{\text{model}}}$), where $d_{\text{model}}=512$ and $c=10$ (context window size), processed through a dropout layer with rate 0.2.
  • Transformer Encoder: 6 layers with 8 attention heads per layer.
  • Feed-Forward Networks: Dimension expansion to 2048 before projection back to $d_{\text{model}}$.
  • Output Layer: Final layer normalization followed by linear projection to $|\mathcal{V}|+1$ dimensions.
Hyperparameter Configuration.
We conducted extensive hyperparameter tuning to optimize our watermarking system’s performance. Table 3 presents our final hyperparameter configuration. For model-specific parameters, we use the OpenCoder-1.5B-Instruct base model with BF16 precision and flash attention 2 implementation for optimal performance. The watermark policy network uses a context window of 2 tokens to determine watermarking decisions, which balances detection effectiveness and computational efficiency. The watermark strength parameters include a green list ratio $\gamma=0.5$, establishing an equal distribution between green and red token sets. The watermark binary decision threshold is set at 0.5, providing a balanced approach for selective watermarking. For detection, we employ a z-score threshold of 4.0, which ensures high precision in distinguishing watermarked code from non-watermarked code while maintaining a low false-positive rate.
C.4 Training Details
We implement reinforcement learning training using GRPO with several customizations for the code watermarking task. Our training pipeline processes the mix of open-r1/verifiable-coding-problems-python dataset, HumanEval dataset and MBPP dataset. The motivation is to make sure model be familiar the distribution of these dataset instructions, which are representative code examples for code generation tasks.
The watermark policy model optimization uses a learning rate of $1.0\times 10^{-5}$ with a cosine scheduler including a minimum learning rate rate of 0.1. We implement a short warmup period (3% of total steps) to stabilize early training. For regularization, we set $\beta=0.0$ for the KL divergence term, allowing the watermark policy to explore freely while process-based rewards provide necessary constraints. The training process runs for 500 steps with a batch size of 8 per device and gradient accumulation every 4 steps. For each prompt, we generate 8 different code completions to provide robust training signals. The maximum prompt length is limited to 256 tokens, while completions can extend up to 1024 tokens to accommodate various code generation tasks. During inference, we use a temperature of 1.0 to maintain natural generation characteristics.
Computation Resources
All experiments were conducted on 1 Nvidia A100 GPU with 80GB memory each, using BF16 precision to maximize computational efficiency. The watermark policy network training requires significantly fewer resources than full LLM training. For inference and evaluation, a single A100 GPU is sufficient to process the benchmark test cases. The lightweight nature of our watermark policy ensures minimal computational overhead during both training and deployment compared to LLM-based watermarking approaches. Our implementation uses PyTorch with flash attention 2 for optimized transformer operations. To ensure reproducibility, we set a random seed of 42 across all experiments. The relatively modest computational requirements make our approach practical for real-world deployment scenarios, as the watermark policy can be integrated with various LLMs without substantial infrastructure demands.
Appendix D Security Discussion
The security of CodeTracer’s watermarking framework derives from the complexity of our learning-based approach and the semantic awareness of our watermark model. This section analyzes the security properties of our approach compared to traditional watermarking methods.
Our transformer-based watermark model provides strong security through its complex neural representations. Unlike static watermarking schemes that apply fixed rules, our architecture learns context-dependent watermarking decisions that adapt to the syntactic and semantic structure of code. The high-dimensional parameter space of our 6-layer transformer model creates a complex mapping between code contexts and watermarking decisions that resists reverse engineering. This neural complexity means that even with access to watermarked outputs, attackers face significant computational barriers to inferring the underlying decision boundaries that govern watermark placement and token partitioning.
The selective application of watermarking through the binary decision mechanism (w) adds another layer of security. Our model strategically applies watermarking only in positions where it can effectively bias toward green tokens without compromising code functionality. This selective approach creates an irregular watermarking pattern that varies with code structure and semantics, making statistical detection more challenging for attackers attempting to identify and remove watermarks.
In contrast, traditional watermarking approaches relying solely on hash functions or fixed vocabulary partitioning remain vulnerable to statistical attacks. Attackers can analyze token distribution patterns across multiple watermarked samples to eventually infer the underlying partitioning scheme. Our framework avoids this vulnerability by creating position-specific, semantically-informed token partitions that vary based on code context.
For a successful attack, adversaries would need to simultaneously: (1) reverse-engineer the neural network’s complex decision boundaries, (2) understand the dynamic token partitioning strategy across different code structures, and (3) develop a method to generate valid code that avoids watermarked patterns without compromising functionality. This multi-faceted security approach provides robust protection against both detection evasion and watermark forgery in practical deployment scenarios.

### v2

@@
-Optimizing Token Choice for Code Watermarking: A RL Approach
+Optimizing Token Choice for Code Watermarking: An RL Approach
 Abstract
-The need for detecting LLM-generated code necessitates watermarking systems capable of operating within its highly structured and syntactically constrained environment.
-To address this, we introduce CodeTracer, an innovative adaptive code watermarking framework underpinned by a novel reinforcement learning training paradigm.
+Protecting intellectual property on LLM-generated code necessitates effective watermarking systems that can operate within code’s highly structured, syntactically constrained nature.
+In this work, we introduce CodeTracer, an innovative adaptive code watermarking framework underpinned by a novel reinforcement learning training paradigm.
 At its core, CodeTracer features a policy-driven approach that utilizes a parameterized model to intelligently bias token choices during next-token prediction.
@@
 Extensive comparative evaluations demonstrate CodeTracer’s significant superiority over state-of-the-art baselines in both watermark detectability and the preservation of generated code’s functionality.
+Our code is available at https://anonymous.4open.science/r/CodeTracer-B8EE.
 1 Introduction
-The unprecedented capabilities of large language models (LLMs) in code generation have introduced critical challenges for intellectual property protection and code attribution [27, 1].
-As AI systems produce increasingly sophisticated code that is difficult to distinguish from human-written, the need for reliable code tracing approaches has become urgent [26].
-Code watermarking, a technique historically vital for software protection, now serves as the cornerstone for identifying and verifying machine-generated code origins [9, 11, 4].
-Unlike natural language, code presents unique watermarking challenges due to its strict structural and semantic requirements: even slight modifications can completely break functionality, a problem less prevalent in text watermarking [13].
-LLM generation is an autoregressive process composed of sequential next-token predictions [20].
-LLM watermarking biases these predictions toward token choices that subtly differ from normal selections, creating detectable statistical patterns.
-Many LLM watermarking schemes typically rely on a “green-red list” methodology [9].
-This technique biases token generation toward specific green (preferred) tokens and suppressing red (avoided) tokens, creating detectable statistical patterns without significantly degrading output quality.
-For natural language, these vocabulary partitions can be determined randomly [9, 13] with minimal impact on coherence.
-Code watermarking faces unique challenges: random token selection often breaks functionality since certain tokens are syntactically essential, while uniform watermarking ignores structural constraints where some positions cannot tolerate modifications.
-Therefore, we identify our core challenge: how to intelligently determine optimal watermark insertion points and select semantically reasonable token choices that maintaining statistical detectability patterns while preserve code functionality?
-Traditional code watermarking approaches suffer from significant limitations: they operate as post-generation transformations on completed code, making them incompatible with the causal generation process of LLMs.
-Additionally, these methods require labor-intensive, language-specific transformation rules that must be manually crafted for each programming language [8, 24, 14].
-Recent efforts to incorporate watermarks during LLM code generation show promise but remain impractical in real-world scenarios, as they require access to supplementary prompts and model information to compute critical values like entropy during detection [13, 6].
+The unprecedented capabilities of large language models (LLMs) in code generation have introduced critical challenges for intellectual property protection and code attribution (Li et al., 2022; Achiam et al., 2023; Guo et al., 2024; Hui et al., 2024).
+As AI systems produce increasingly sophisticated code that is difficult to distinguish from human-written code, the need for reliable code tracing approaches has become urgent (Zhao et al., 2024; Wang et al., 2024).
+Traditional code watermarking approaches apply post-generation transformations to completed code, making them inherently incompatible with the autoregressive generation process of LLMs.
+Moreover, these methods require labor-intensive, language-specific transformation rules that must be manually crafted for each programming language (Hamilton & Danicic, 2011; Yang et al., 2023; Li et al., 2024).
+Existing LLM watermarking approaches operate on the autoregressive generation process by biasing next-token predictions toward statistically detectable patterns (Radford et al., 2019; Kirchenbauer et al., 2023a).
+In natural language contexts, this approach succeeds because text generation is robust, as most positions admit multiple semantically valid token choices, allowing watermarking to be applied across all generated tokens (Kuditipudi et al., 2023; Dathathri et al., 2024; Liu & Bu, 2024).
+However, code watermarking presents distinct challenges that arise from the structural and semantic constraints inherent to programming languages.
+Code generation imposes two critical constraints.
+First, syntactic dependencies severely constrain the space of valid token choices, as certain tokens are syntactically mandatory and their modification results in compilation failures (Guan et al., 2024).
+Second, code positions exhibit heterogeneous sensitivity to modifications, where indiscriminate watermarking strategies fail to account for the varying tolerance to perturbations across different code locations (Lee et al., 2023).
+Recent efforts to incorporate watermarks during LLM code generation show promise but remain impractical in real-world scenarios, as they require access to supplementary prompts and model information to compute critical values like entropy during detection (Lee et al., 2023; Guan et al., 2024; Zhao et al., 2024).
 We argue that an effective solution requires a model-based approach that learns the necessary programming knowledge during training, enabling intelligent watermarking decisions that adapt to syntactic and semantic constraints with minimal context.
-In this work, we propose CodeTracer, an adaptive code watermarking framework built on a novel reinforcement learning (RL) paradigm.
-Our approach features a parameterized watermark model that intelligently identifies optimal positions and biases token selection during the code generation process.
-Reinforcement learning provides an ideal solution for our task as it eliminates the need for pre-existing watermarked code examples.
-This is crucial because creating such training examples presents a fundamental paradox: we cannot determine whether a code snippet is “watermarked” without first defining the watermark model itself.
-Challenges with RL.
-Designing a learnable code watermarking RL framework presents significant challenges.
-Two critical difficulties emerge in this process.
-First, effective reward mechanisms are fundamental to guiding the RL process, requiring careful integration of available feedback signals.
-Second, the watermarking procedure necessitates making discrete decisions about code tokens, creating a non-differentiable process that is inherently difficult to optimize.
-To address these challenges, we implement a dual-component reward system: execution feedback that penalizes functionally incorrect code, complemented by watermark embedding signals consisting of both immediate process rewards for successful watermarked token selection and statistical outcome rewards that employ metrics like z-score to comprehensively measure detectability.
-Additionally, we overcome the non-differentiability constraint through Gumbel Top-k reparameterization [22], enabling gradient-based optimization by allowing differentiable selection of tokens.
+Consequently, we identify the core challenge in LLM code watermarking: how can we intelligently determine optimal watermark insertion points and select semantically reasonable token choices that maintain statistical detectability while preserving code functionality?
+In this work, we propose CodeTracer, an adaptive code watermarking framework built upon a novel reinforcement learning (RL) training paradigm.
+CodeTracer employs a policy-driven approach that leverages a parameterized model to intelligently identify optimal insertion positions and guide token selection during the code generation process.
+The parameterized model collaborates with the original LLM to form a watermarked policy, where only the parameterized model undergoes optimization during training.
+At generation time, the LLM provides logits while the parameterized model intelligently determines watermark application and token selection decisions.
+To train the watermarked policy, we employ reinforcement learning through a carefully designed dual-component reward system.
+This system integrates two complementary feedback signals: execution feedback that penalizes functionally incorrect code, and watermark embedding signals that comprise both immediate process rewards for successful watermarked token selection and statistical outcome rewards employing metrics such as the z-score to comprehensively assess detectability performance.
+To enable end-to-end gradient-based optimization, we address the non-differentiability of discrete bias token selection by Gumbel Top-k reparameterization (Xie & Ermon, 2019) and Straight-Through Estimation (Bengio et al., 2013), achieving differentiable watermarked policy training.
 Contributions.
-Our work makes several key contributions: (i) We introduce CodeTracer, an adaptive watermarking framework that intelligently embeds watermarks during LLM code generation; (ii) We develop a novel RL pipeline for code watermarking, combines execution feedback with dual watermark signals, and enables differentiable optimization for the whole pipeline; and (iii) Our evaluations across multiple datasets demonstrate CodeTracer’s superior watermark detection capabilities while maintaining code functionality.
-2 Related Works
+Our work yields several key contributions: (i) We introduce CodeTracer, an adaptive watermarking framework that intelligently embeds watermarks during LLM code generation; (ii) We develop a novel RL pipeline for code watermarking training that combines execution feedback with dual watermark signals and enables differentiable optimization for the entire pipeline; and (iii) Empirically, we validate the effectiveness of CodeTracer through comprehensive evaluations, demonstrating its superior watermark detection capabilities while maintaining code functionality.
+2 Related Work
 LLM Watermarking.
-LLMs’ sophisticated content generation capabilities create challenges for authenticity verification [26, 19, 4].
-Recent watermarking approaches embed signatures during the token sampling process by modifying logits or altering the sampling procedure [9, 11].
-A notable example is the green-red watermarking scheme [9], which partitions vocabulary tokens into lists and biases selection toward “green” tokens, enabling statistical detection.
-[23] introduces a RL-based approach for watermark embedding but their co-training scheme may cause unexpected behavior for LLMs.
-However, these methods struggle with low-entropy conditions common in code generation [13].
+Existing LLM watermarking approaches embed imperceptible signatures during token sampling by modifying logits or altering the sampling procedure (Kirchenbauer et al., 2023a; Kuditipudi et al., 2023; Zhao et al., 2023; Christ et al., 2024; Dathathri et al., 2024).
+A prominent example is the green-red watermarking scheme (Kirchenbauer et al., 2023a; b), which partitions the vocabulary into “green” (preferred) and “red” (avoided) tokens, biasing generation toward green tokens while suppressing red ones, enabling statistical detection of watermarked content.
+Xu et al.
+(2024) propose a reinforcement learning approach for watermark embedding that requires training the LLM, which may cause unexpected behaviors for LLMs.
+Critically, these methods encounter difficulties in low-entropy scenarios typical of code generation (Lee et al., 2023).
 Code Watermarking.
-Code watermarking imposes unique challenges due to programming languages’ strict syntactic and semantic requirements.
-Traditional code watermarking alters existing code through format changes or control flow restructuring [8, 17, 14, 24, 16, 4].
-Recent LLM-focused approaches use entropy distributions [13, 15] or type predictors [6] to guide watermark insertion.
-However, these techniques typically require privileged access to LLM parameters, generation probabilities, or original prompts, limiting their practical deployment.
-CodeTracer instead adaptively embeds watermarks during generation without such prerequisites.
-Rule-Based RL and GRPO.
-Rule-Based Reinforcement Learning enhances LLMs by integrating explicit rules into the learning process, demonstrating reduced vulnerability to reward hacking [7, 12].
-These approaches leverage simple, verifiable rules that ground model outputs, showing particular effectiveness on structured tasks like coding and mathematical reasoning [25].
-Deepseek-R1 [7] exemplifies this approach through Group Relative Policy Optimization (GRPO), which significantly reduces computational overhead while improving complex reasoning capabilities.
-GRPO generates multiple candidate responses for a given prompt, then calculates advantages by normalizing rewards relative to group statistics, eliminating the need for trained critic models in value estimation.
-Our CodeTracer framework adopts this GRPO scheme with rule-based rewards for computational efficiency.
-3 Preliminaries
-LLM Watermarking.
-In a code generation task, large language model receives a sequence of tokens $\mathbf{x}=[x_{1},x_{2},\ldots,x_{n}]$ as the input prompt, and generates output $\mathbf{y}=[y_{1},y_{2},\ldots,y_{m}]$ including the code snippet.
-At every generation step $t$, the LLM $\pi_{\theta}$ receives the input prompt $\mathbf{x}$ and previous generated token to compute the logit vector $\mathbf{l}=[l_{1},l_{2},\ldots,l_{|\mathcal{V}|}]$ for token $t$, cross all tokens in the vocabulary $\mathcal{V}$, where each $l_{j}$ corresponds to token $v_{j}\in\mathcal{V}$..
-The logit vector is then applied with a softmax opreator to transform it into a probability distribution where token at position $y_{t}$ is sampled from.
-To detect AI generated content, in-generation watermarking [9] adjust the original LLM generation policy by adding constant biases on a sublist of vocabulary that are selected by a pseudorandom function, resulting in a watermarked code sequence $\mathbf{\tilde{y}}=[\tilde{y}_{1},\tilde{y}_{2},\ldots,\tilde{y}_{m^{\prime}}]$.
-The statistical difference in the token distribution within/out the sublists are then utlized to detect the AI-generated output.
-Challenge of LLM Code Watermarking.
-However, LLM watermark faces serve performance degration in the code because of unique challenges distinct from natural text watermarking.
-Unlike natural language, code is highly structured with a delicate syntax where small modifications can drastically alter functionality or render the code inoperable.
+Code watermarking presents distinct challenges due to the strict syntactic and semantic constraints inherent in programming languages.
+Traditional approaches modify existing code through formatting changes or control flow restructuring (Hamilton & Danicic, 2011; Ma et al., 2019; Li et al., 2024; Yang et al., 2023; Liu et al., 2024; Dathathri et al., 2024).
+Recent LLM-focused methods leverage entropy distributions (Lee et al., 2023; Li et al., 2023) or type predictors (Guan et al., 2024) to guide watermark insertion.
+However, these techniques typically require privileged access to LLM parameters, generation probabilities, or original prompts during watermark detection, significantly limiting their practical deployment (Zhao et al., 2024).
+In contrast, CodeTracer adaptively embeds watermarks during generation without such prerequisites.
+RLVR and GRPO.
+Reinforcement Learning with Verifiable Rewards (RLVR) enhances LLMs by integrating verifiable reward signals into the training process, demonstrating improved robustness against reward hacking and superior performance (Lambert et al., 2024; Guo et al., 2025; Team et al., 2025).
+However, RLVR is constrained by scarcity of reliably verifiable signals, with most applications limited to mathematical problems and code execution tasks.
+DeepSeek-R1 (Guo et al., 2025) combines verifiable rewards with Group Relative Policy Optimization (GRPO), which improves computational efficiency over traditional policy optimization methods.
+Code watermarking emerges as a natural fit for RLVR, as watermark detection provides unambiguous, token-level verification signals that can be efficiently computed.
+We exploit this natural alignment by employing GRPO with watermark-based verifiable rewards to achieve both computational efficiency and robust watermarking.
+3 CodeTracer: A Policy-Driven Watermarking Framework
+Problem Setting.
+In code generation tasks, a large language model takes a sequence of input tokens representing the prompt $\mathbf{x}=[x_{1},x_{2},\ldots,x_{n}]$ and generates an output sequence $\mathbf{y}=[y_{1},y_{2},\ldots,y_{m}]$ containing the generated code.
+At each generation step $t$, the LLM $\pi_{\theta}$ processes the input prompt $\mathbf{x}$ and previously generated tokens to compute a logit vector $\mathbf{l}=[l_{1},l_{2},\ldots,l_{|\mathcal{V}|}]$ over the entire vocabulary $\mathcal{V}$.
+Each value $l_{j}$ represents the model’s preference for token $v_{j}\in\mathcal{V}$.
+This logit vector is transformed into a probability distribution using the softmax function, from which token $y_{t}$ is sampled.
+To enable detection of LLM-generated code, in-generation watermarking techniques, such as the approach by Kirchenbauer et al.
+(2023a), modify the LLM’s original logits $\mathbf{l}$ to form watermarked logits $\mathbf{\tilde{l}}$.
+This modification is achieved by adding biases to the logits of a specific subset (green list) $G\subset\mathcal{V}$ of the vocabulary, selected using a pseudorandom function (PRF).
+This process yields a watermarked code sequence $\mathbf{\tilde{y}}=[\tilde{y}_{1},\tilde{y}_{2},\ldots,\tilde{y}_{m^{\prime}}]$.
+The presence of watermark is then statistically inferred by analyzing frequency difference of tokens appearing within the biased vocabulary subset.
+Challenges in LLM Code Watermarking.
+LLM watermarking faces severe performance degradation in code generation due to unique challenges distinct from natural text watermarking.
+Unlike natural language, code is highly structured with precise syntax where small modifications can drastically alter functionality or render the code inoperable.
 This inherent rigidity creates significant constraints for watermarking techniques.
-First, watermark position selection is particularly critical in code, as certain positions are immutable.
-Unlike natural text where minor alterations rarely affect meaning, code contains strict structural elements that cannot be modified.
-For example, altering the def keyword in a Python function definition (e.g., changing def process_data(data): to func process_data(data):) inherently breaks syntax.
-Any watermarking approach must identify and avoid these critical positions to maintain functional code.
+First, watermark position selection is critical, as many positions in code are immutable.
+Unlike natural text where alterations rarely affect meaning, code contains structural elements that cannot be modified.
+For example, changing def to func in Python function definitions breaks syntax.
+Any watermarking approach must avoid these critical positions.
 Second, watermark token choice must respect contextual constraints.
-Even at positions where modification is possible, the replacement tokens must be semantically and syntactically appropriate.
-For instance, if a model generates an assignment like status = "active" but a watermarking scheme forces a substitution of "active" with an incompatible token such as the keyword class, the resultant status = class would cause a syntax error.
-This contextual sensitivity significantly limits the vocabulary available for watermarking while preserving code functionality.
-Why RL for $\pi_{\phi}(a|\mathbf{c})$ training?
-Reinforcement learning provides an ideal framework for this task for three key reasons: (i) Learning complex rules implicitly: The diversity of programming languages makes manually defining watermarking rules impractical.
-RL can learn syntactic constraints across various code structures through interaction rather than exhaustive specification.
-(ii) Overcoming the absence of supervised data: SFT is impossible due to a circular dependency—we cannot create "gold standard" watermarked examples without first defining the watermark model itself.
-RL enables learning optimal strategies without requiring predefined examples.
-(iii) Leveraging clear feedback signals: Both watermark effectiveness and code functionality provide quantifiable rewards.
-Our system utilizes rule-based feedback mechanisms that have proven effective for code tasks [7, 12], making RL well-suited for optimizing within code generation constraints.
-4 Learn to Watermark
-We formulate the LLM code watermarking problem within a reinforcement learning (RL) framework.
-This section first introduces the watermark policy, then outlines the process of training this policy through RL, and finally addresses specific design challenges along with our proposed solutions.
-4.1 Watermarked Policy
-We focus on a single step in the token generation process, specifically one next token prediction.
-Watermark model $\pi_{\phi}(a|\mathbf{c})$ operates based on the current context $\mathbf{c}$ (comprising part of input $\mathbf{x}$ and previously generated tokens in $\mathbf{y}$ up to the current position with a fixed number of tokens).
-This model outputs an action $a=(w,G)$ consisting of two components: a binary decision $w\in\{0,1\}$ indicating whether to apply watermarking at this position, and a green token list $G\subset\mathcal{V}$ for preferred tokens.
-The watermarked token generation process combines the original LLM $\pi_{\theta}$ with the watermark model $\pi_{\phi}$ and get watermarked policy $\pi_{\theta\oplus\phi}$, computing the modified logits as:
-$\tilde{l}_{j}=l_{j}+w\cdot\delta\cdot\mathbbm{1}_{v_{j}\in G}$ (1)
-where $\delta$ represents a bias added exclusively to green tokens when watermarking is active ($w=1$), and $\mathbbm{1}_{v_{j}\in G}$ is an indicator function equaling 1 if token $v_{j}\in G$ and 0 otherwise.
-The watermark strength is controlled by two constraints: the size of the green token set where $|G|=\gamma|\mathcal{V}|$ for a predefined parameter $\gamma\in(0,1)$, and the bias value $\delta$.
-The remaining tokens form a “red” group $R=\mathcal{V}\setminus G$.
-This policy-driven, selective approach dynamically controls both watermark placement and vocabulary partitioning used for biasing.
-With this formulation, the policy strategically biases token choices through the green set $G$ and controls watermark application via the binary decision $w$.
-The core challenge lies in training the parameters $\phi$ of the watermark policy model.
-In the following subsection, we present the complete watermark pipeline for both watermark generation and detection using the combined policy $\pi_{\theta\oplus\phi}$, followed by our training methodology that optimizes the watermark policy $\pi_{\phi}$ in next section.
-Generation
-During code generation, our framework CodeTracer applies the previously defined framework sequentially.
-The watermarked policy $\pi_{\theta\oplus\phi}(\mathbf{\tilde{y}}|\mathbf{x})$ takes exactly same input with LLM $\pi_{\theta}$ but generate the biased output sequences $\mathbf{\tilde{y}}=[\tilde{y}_{1},\tilde{y}_{2},\ldots]$ based on biased logits $\tilde{l}_{j}$ at each step.
+Even at modifiable positions, replacement tokens must maintain syntactic validity.
+For instance, if status = ‘‘active’’ is watermarked by replacing ‘‘active’’ with class, the result causes a syntax error.
+This contextual sensitivity severely limits the available vocabulary for watermarking.
+Consequently, effective code watermarking requires a deeper understanding of code structure and semantics.
+3.1 CodeTracer
+To address the aforementioned challenges, CodeTracer introduces a policy-driven framework that integrates a watermark model $\pi_{\phi}$ with an LLM $\pi_{\theta}$, yielding a composite watermarked policy $\pi_{\theta\oplus\phi}$ capable of generating watermarked code.
+At each generation step $t$, the watermark model $\pi_{\phi}(a|\mathbf{c})$ operates conditioned on the current context $\mathbf{c}$, defined as the concatenation of a segment of the input $\mathbf{x}$ and the sequence of previously generated tokens $\mathbf{y}_{<t}$ within a fixed-length window.
+The output of this policy is an action $a=(w,G)$, where $w\in\{0,1\}$ is a binary variable indicating whether to apply watermarking at the current position, and $G\subset\mathcal{V}$ represents a set of preferred “green” tokens.
+The generation of a token is then achieved by sampling from a modified logit vector from the watermarked policy $\pi_{\theta\oplus\phi}$ as:
+$\tilde{l}_{j}=l_{j}+w\cdot\delta\cdot\mathbbm{1}_{v_{j}\in G},$ (1)
+where $\delta$ is a hyperparameter controlling the bias applied to the logits of tokens in the green list $G$ when watermarking is active ($w=1$), and $\mathbbm{1}_{v_{j}\in G}$ is an indicator function equaling 1 if token $v_{j}\in G$ and 0 otherwise.
+The watermark strength is governed by the size of the green token set, $|G|=\gamma|\mathcal{V}|$ for a predefined ratio $\gamma\in(0,1)$, and the bias magnitude $\delta$.
+The complement of the green set, $R=\mathcal{V}\setminus G$ constitutes the “red” token set.
+This policy-driven and context-aware approach allows for dynamic control over both the placement of the watermark and the vocabulary subset used for biasing the generation process.
+In essence, this formulation empowers the policy to strategically influence the token sampling process by preferentially selecting tokens from the green set $G$, while simultaneously regulating the watermark through the binary decision $w$.
+Consequently, during generation, CodeTracer applies the watermarked policy $\pi_{\theta\oplus\phi}(\mathbf{\tilde{y}}|\mathbf{x})$ that takes the same input as the LLM $\pi_{\theta}$ but generates biased outputs $\mathbf{\tilde{y}}=[\tilde{y}_{1},\tilde{y}_{2},\ldots]$ via modified logits $\tilde{l}_{j}$.
 Detection.
-For detection, CodeTracer analyzes a potentially corrupted sequence $\mathbf{s}=[s_{1},s_{2},\ldots,s_{T^{\prime}}]$ of length $T^{\prime}$.
-The watermark policy $\pi_{\phi}$ reconstructs its decisions independently, without requiring the LLM $\pi_{\theta}$.
-This reconstruction process determines two critical pieces of information for each token: first, whether watermarking was applied at that position ($w=1$ or $w=0$), and second, which tokens belonged to the green list $G$ if watermarking was applied.
-By applying this reconstruction, we identify the subset of tokens $s_{w=1}$ where watermarking was active, yielding a total count of watermarked positions $T=|\{s_{w=1}\}|$.
-Among these watermarked positions, we then count how many tokens fall into their respective green sets, denoted as $N_{G}=|\{s_{w=1,s\in G}\}|$.
-A one-proportion $z$-test assesses statistical significance:
+During detection, given an output sequence $\mathbf{s}=[s_{1},s_{2},\ldots,s_{T^{\prime}}]$ of length $T^{\prime}$, CodeTracer reconstructs the watermarking decisions independently using only the watermark model $\pi_{\phi}$, without requiring access to the LLM $\pi_{\theta}$.
+This reconstruction determines, for each token, whether watermarking was applied ($w=1$ or $w=0$) and, if so, the composition of the corresponding green token set $G$.
+We identify the subset of tokens where watermarking was active, denoted as $\{s_{w=1}\}$, and count the total number of watermarked positions $T=|\{s_{w=1}\}|$.
+We then count how many of watermarked tokens belong to their respective reconstructed green sets, denoted as $N_{G}=|\{s\mathrel{\mathop{\ordinarycolon}}s\in\{s_{w=1}\}\land s\in G\}|$.
+To assess the statistical significance, we employ a one-proportion $z$-test for the proportion of watermarked tokens belonging to their predicted green sets:
 $z=\frac{N_{G}-T\gamma}{\sqrt{T\gamma(1-\gamma)}}.$ (2)
-A significantly large positive $z$-score indicates watermark presence, demonstrating that watermarked tokens appear in their respective green sets more frequently than would occur by random chance.
+Under the null hypothesis of no watermarking, we expect $N_{G}$ to follow a binomial distribution with success probability $\gamma$.
+A sufficiently large positive $z$-score provides strong statistical evidence for watermark presence, indicating that watermarked tokens appear in their predicted green sets with frequency significantly higher than the expected random baseline of $\gamma$.
+The complete algorithms for watermark generation and detection are formalized in Algorithm 1 and Algorithm 2, respectively.
 Figure 1: CodeTracer: A framework for LLM code watermarking through selective token biasing.
-The diagram illustrates the end-to-end pipeline where a trainable watermark model works alongside an LLM to embed statistical patterns in code, with a reward system that optimizes for both code functionality and watermark detectability.
+The diagram shows our end-to-end pipeline where a trainable watermark model collaborates with an LLM to embed detectable statistical patterns in generated code.
+A reward system optimizes the dual objectives of preserving code functionality while maximizing watermark detectability.
+The watermark model operates as a plug-in module, enabling deployment beyond those used during training.
+Importantly, watermark detection requires only the watermark model, not the original LLM.
+4 Learning to Watermark in CodeTracer
+We formulate training of watermarked policy $\pi_{\theta\oplus\phi}$ within a reinforcement learning framework.
+This section first gives an overview of the training pipeline, then discusses the reason to use RL for policy learning, and finally addresses specific design challenges along with our proposed approach.
 4.1 How to Learn the Watermarked Policy?
-As shown in Figure 1, to train the watermark model $\pi_{\phi}(a|\mathbf{c})$ in the watermarked policy $\pi_{\theta\oplus\phi}$, we employ GRPO as our RL framework.
-We selected GRPO for its computational efficiency and demonstrated effectiveness when integrating rule-based rewards on coding and mathematical tasks [7].
-In a standard GRPO training scheme, three components are essential: a training policy, a reference policy, and reward functions.
-Our training pipeline uses $\pi_{\theta\oplus\phi}$ as the training policy rather than $\pi_{\phi}$ alone, since the complete generation process requires the coordinated operation of both the LLM and the watermark model.
-Importantly, we freeze the LLM parameters $\theta$ during training and optimize only watermark model parameters $\phi$, ensuring a portable and efficient watermark solution.
+As shown in Figure 1, we employ the GRPO algorithm (Shao et al., 2024) as our RL framework to train the watermark model $\pi_{\phi}(a|\mathbf{c})$ within the watermarked policy $\pi_{\theta\oplus\phi}$.
+GRPO offers computational efficiency and has demonstrated effectiveness in integrating rule-based rewards for coding and mathematical tasks (Guo et al., 2025).
+Our training pipeline requires three key components from the standard GRPO framework: a training policy, a reference policy, and reward functions.
+We use the integrated watermarked policy $\pi_{\theta\oplus\phi}$ as the training policy rather than solely $\pi_{\phi}$, as the complete code generation process requires coordinated operation of both the LLM and watermark model.
+Crucially, we freeze the LLM parameters $\theta$ during training and optimize only the watermark model parameters $\phi$.
+This design ensures a portable and efficient watermarking solution that trained watermark model can be applied to other pre-trained LLMs without requiring fine-tuning or additional training data.
+4.2 Why Use RL for Learning?
+Code watermarking presents a fundamental training paradox that reinforcement learning uniquely addresses.
+Traditional supervised learning requires pre-existing watermarked examples to train a detector, but creating such examples necessitates an already-trained watermark model: a circular dependency that renders supervised approaches infeasible.
+Reinforcement learning eliminates this paradox by learning through interaction rather than predefined examples.
+Code watermarking is naturally suited for RL due to two key characteristics.
+First, watermark detection provides unambiguous, token-level verification signals that can be efficiently computed.
+Second, code functionality offers clear binary feedback through test case execution.
+Our RL formulation leverages these signals to offer distinct advantages: (1) it implicitly learns complex syntactic constraints across diverse programming languages without exhaustive manual specification; (2) it optimizes the watermarked policy to balance detectability and functionality through quantifiable reward mechanisms; (3) it enables end-to-end training without requiring pre-watermarked data.
 4.3 Policy Design
-The implementation of our watermarked policy $\pi_{\theta\oplus\phi}$ presents fundamental technical challenges due to the non-differentiable nature of several components in Equation (1).
-Specifically, we implement the watermark model $\pi_{\phi}(a|\mathbf{c})$ as a transformer-based architecture.
-While the transformer outputs are continuous, we need to convert them to discrete outputs while maintaining differentiability.
-The watermark binary decision output $w$ and green token list assignment $G$ involve discrete selection processes that create optimization barriers in the neural network training pipeline.
+The implementation of our watermarked policy $\pi_{\theta\oplus\phi}$ presents fundamental technical challenges due to the discrete nature of the watermark components $w$ and $G$ in Equation equation 1.
+We implement the watermark model $\pi_{\phi}(a|\mathbf{c})$ where $a=(w,G)$ as a transformer that outputs continuous values, but we need discrete outputs for watermarking decisions.
+Both the binary watermark decision $w$ and the green token list $G$ require discrete selections that impede gradient flow during training.
 Straight-Through Estimation.
-To enable gradient flow through the discrete decision processes for $w$, we apply straight-through estimation techniques.
-The watermark model transformer architecture first outputs a $|l_{\phi}|+1$ dimensional vector, which we represent as $(w_{\phi},l_{\phi})$.
-To derive our desired output $a=(w,G)$ from $(w_{\phi},l_{\phi})$, we implement the straight-through estimations:
-$w=\mathbbm{1}_{w_{\phi}>0.5}+\sigma(w_{\phi})-\text{sg}(\sigma(w_{\phi})).$ (3)
-During the forward pass, this formulation behave like the discrete operations $\mathbbm{1}[w_{\phi}>0]$.
-However, during backpropagation, gradients flow through the continuous relaxations $\sigma(w_{\phi})$, with $\text{sg}(\cdot)$ ensuring that only the gradient of the relaxation (and not their forward values) affects the computation.
-This technique enables end-to-end training despite the discrete nature of watermarking decisions.
+To enable gradient flow through the discrete decision processes for watermark selection $w$, we apply straight-through estimation techniques.
+Specifically, our watermark transformer outputs a $(|\mathcal{V}|+1)$-dimensional vector as $(w_{\phi},l_{\phi})$, where $w_{\phi}\in\mathbb{R}$ is a scalar for watermark placement probability of and $l_{\phi}\in\mathbb{R}^{|\mathcal{V}|}$ as logit bias vector.
+To derive our desired output $a=(w,G)$ from $(w_{\phi},l_{\phi})$, we implement straight-through estimation:
+$w=\mathbbm{1}_{w_{\phi}>0}+\sigma(w_{\phi})-\text{sg}(\sigma(w_{\phi})),$ (3)
+where $\sigma(\cdot)$ is the sigmoid function and $\text{sg}(\cdot)$ is the stop-gradient operator.
+During the forward pass, this formulation behaves like the discrete operation $\mathbbm{1}_{w_{\phi}>0}$.
+However, during backpropagation, gradients flow through the continuous relaxation $\sigma(w_{\phi})$.
+The stop-gradient operator ensures only the gradient of the relaxation (and not its forward values) affects the computation.
+This technique enables end-to-end training despite the discrete nature of watermarking decisions made by $w$.
 Gumbel-Top-k.
@@
 Given logits $\mathbf{l}_{\phi}$ from the watermark model $\pi_{\phi}$, we apply Gumbel noise to create a continuous relaxation of the discrete selection:
-$\mathbf{g}=\mathbf{l}_{\phi}+(-\log(-\log(\mathbf{u})))$ (4)
+$\mathbf{g}=\mathbf{l}_{\phi}+(-\log(-\log(\mathbf{u}))),$ (4)
 where $\mathbf{u}\sim\text{Uniform}(0,1)^{|\mathcal{V}|}$ represents uniform random noise across the vocabulary space.
 The green list $G$ is then determined by selecting the top-$k$ tokens according to these perturbed logits:
-$G=\text{arg top-{k}}(\mathbf{g})$ (5)
+$G=\text{arg top-{k}}(\mathbf{g}).$ (5)
 To enable gradient flow through this discrete selection process, we implement a straight-through estimator for the indicator variable $\mathbf{l}_{G}\in\{0,1\}^{|\mathcal{V}|}$ that represents membership in the green list:
-$\mathbf{l}_{G}=\mathbbm{1}_{v\in G}+\mathcal{S}(\mathbf{g})_{v}-\text{sg}(\mathcal{S}(\mathbf{g})_{v})$ (6)
-where $\mathcal{S}(\mathbf{g})_{v}$ represents the Gumbel-Softmax-based continuous relaxation of the top-$k$ selection for token $v$, and $\text{sg}(\cdot)$ is the stop-gradient operator.
-During forward passes, we use hard discrete token selection based on token rankings from the Gumbel-perturbed logits, while during backward passes, the gradients flow through the softmax-based continuous relaxation.
-This dual-path approach allows the model to learn effective watermarking strategies while maintaining the discrete token selection.
-For the reference policy component in our GRPO framework, we utilize the same architecture as the training policy $\pi_{\theta\oplus\phi}$, establishing a self-referential optimization approach.
+$\mathbf{l}_{G}=\mathbbm{1}_{v\in G}+\mathcal{S}(\mathbf{g})-\text{sg}(\mathcal{S}(\mathbf{g})),$ (6)
+where $\mathcal{S}(\mathbf{g})$ represents the Gumbel-Softmax-based continuous relaxation of the top-$k$ selection, and $\text{sg}(\cdot)$ is the stop-gradient operator.
+During forward passes, we use hard discrete token selection $\mathbbm{1}_{v\in G}$ with Equation 5 based on token rankings from the Gumbel-perturbed logits $\mathbf{g}$, while during backward passes, the gradients flow through the softmax-based continuous relaxation $\mathcal{S}(\mathbf{g})$.
+This approach allows the model to learn effective watermarking while maintaining the discrete token selection.
+For the reference policy in our GRPO framework, we utilize the same architecture as the training policy $\pi_{\theta\oplus\phi}$, establishing a self-referential optimization approach.
 This design choice enables the watermarking parameters to be refined against their own previous distributions.
-The approach facilitates gradual refinement of watermark policy while ensuring that successive iterations remain coherent with previous behaviors, ultimately leading to more stable convergence and better performance.
+The approach facilitates gradual refinement of watermarked policy while ensuring that successive iterations remain coherent with previous behaviors, ultimately leading to more stable convergence and better performance.
 4.4 Reward Function Design
-The goal of the reward system is to leverage watermark detection feedback and code execution signals to guide watermarked policy $\pi_{\theta\oplus\phi}$ in selecting suitable watermarking positions ($w$) and making effective token choices for the green list ($G$).
-To achieve this, we have designed a reward system that comprises a set of reward functions balancing the objectives crucial for code watermarking.
-Outcome-based Execution Reward $R_{1}$.
-Our execution reward $R_{1}$ implements a binary evaluation of the watermarked code’s functional integrity through test case performance. We define:
-[…]
-This straightforward formulation provides clear feedback on code correctness: successful execution of all test cases yields a positive reward, while any compilation error or runtime failure incurs a substantial penalty.
-By incorporating this binary signal into our reward function, we establish a strict constraint that compels the watermark policy to preserve the code’s original functionality.
-Outcome-based Watermark Reward $R_{2}$.
-The watermark detection reward $R_{2}$ quantifies the statistical detectability of embedded watermarks within generated code sequences using a saturated function based on the statistical z-score.
-For each complete sequence $s$, we implement:
-$R_{2}(s)=\begin{cases}1,&\text{if }z(s)\geq 3\\ \frac{z(s)}{3},&\text{if }0<z(s)<3\\ 0,&\text{if }z(s)\leq 0\end{cases}$ (8)
-where $z(s)$ represents the statistical z-score from our detection algorithm.
+Our reward system guides the watermarked policy $\pi_{\theta\oplus\phi}$ in watermark placement $w$ and green token selection $G$ using feedback from watermark detection and code execution.
+We design a multifaceted reward system that balances code functionality and watermark detectability.
+Outcome-Based Execution Reward $R_{1}$. Our execution reward $R_{1}$ evaluates code functional correctness through binary assessment based on the performance on passing test cases:
+$R_{1}=\begin{cases}1,&\text{if all test cases pass}\\ 0,&\text{otherwise}\end{cases}$ (7)
+This binary formulation provides unambiguous feedback on code correctness, where successful execution across all test cases yields a positive reward, while any compilation error, runtime failure, or test failure results in zero reward.
+By incorporating this strict binary signal into our reward function, we enforce functional preservation as a hard constraint, ensuring that the watermarked policy maintains code correctness while learning watermarking strategies.
+Outcome-Based Watermark Reward $R_{2}$. The watermark detection reward $R_{2}$ quantifies the statistical detectability of the watermark by employing a saturated function of the statistical $z$-score:
+$R_{2}(s)=\begin{cases}1,&\text{if }z(s)\geq 4\\ \frac{z(s)}{4},&\text{if }0<z(s)<4\\ 0,&\text{if }z(s)\leq 0\end{cases}$ (8)
+where $z(s)$ is the statistical $z$-score obtained from the detection algorithm as a holistic measurement of watermark performance.
 This saturated reward formulation establishes clear boundaries: a $z$-score of 3 or higher (indicating strong statistical significance) receives the maximum reward of 1, while non-positive $z$-scores receive no reward.
 Between these thresholds, the reward scales linearly with the $z$-score.
-This design incentivizes the watermark policy to achieve statistically significant detection levels without wasting computational resources on unnecessarily strong signals.
+This design incentivizes our policy to achieve statistically significant detection levels.
 Process-Based Watermark Reward $R_{3}$.
-In our implementation, we find that outcome-based rewards alone are insufficient for effective policy learning, as they assign equal advantage to both red and green tokens when the overall sequence succeeds.
-To provide more granular guidance, we introduce a process-based reward $R_{3}$ that evaluates each token individually, offering immediate feedback on token selection decisions.
-This approach is straightforward to implement as we can directly check whether each generated token $s_{t}$ is green, red, or non-watermarked based on the policy’s action $a_{t}=(w_{t},G_{t})$ at the $t$-th token:
-[…]
-This simple yet effective reward structure explicitly encourages the selection of green tokens during watermarking while penalizing red token selection.
-Non-watermarked positions receive a neutral reward, focusing the learning signal on positions where watermarking is applied.
-By providing direct feedback at each token generation step, the process-based reward creates a much clearer learning landscape compared to sequence-level outcome rewards alone.
-This token-by-token guidance significantly accelerates policy convergence and leads to more effective watermarking strategies.
-Table 1: Performance (%) comparison of text detection methods.
-Methods marked with work in an impractical scenario with original LLM and prompts.
-Bold values indicate best performance within each category.
+We observed that relying solely on outcome-based rewards can be insufficient for effective policy learning, as they provide the identical reward for tokens in a sequence.
+To provide more granular guidance, we introduce a process-based reward $R_{3}$ that evaluates individual generated tokens, offering immediate feedback on token selection decisions.
+This reward is straightforward to implement, as we can directly assess whether each generated token $s_{t}$ is green, red, or non-watermarked based on the policy’s action $a_{t}=(w_{t},G_{t})$ at time $t$:
+$R_{3}(s_{t},a_{t})=\begin{cases}1,&\text{if }w_{t}=1\text{ and }s_{t}\in G_{t}\\ -1,&\text{if }w_{t}=1\text{ and }s_{t}\in R_{t}\\ 0,&\text{if }w_{t}=0\end{cases}$ (9)
+This reward structure explicitly encourages green token selection during watermarking while penalizing red token selection.
+Non-watermarked positions receive neutral rewards, concentrating the learning signal on watermarked tokens.
+By providing immediate feedback at each generation step, $R_{3}$ delivers more granular guidance than sequence-level rewards alone, significantly accelerating policy convergence and improving watermarking performance through direct token-level supervision.
+Integration into Advantage Computation.
+Our approach calculates advantages at two complementary levels to optimize the watermarked policy.
+First, we compute outcome advantages $A_{1}$ from the execution reward $R_{1}$ and watermark detection reward $R_{2}$ using GRPO’s outcome-based group normalization mechanism.
+Second, we derive process advantages $A_{2}$ from the token-level watermark reward $R_{3}$, normalizing these values across all tokens in all rollouts for each prompt.
+We categorize generated tokens into four distinct types: green tokens (watermarked and belonging to set $G$), red tokens (watermarked but falling in set $R$), non-watermarked tokens (where $w=0$), and non-code text tokens.
+To maintain a specific focus on code watermarking, we exclude non-code tokens entirely.
+The final advantage function for each token $s_{t}$ at time $t$ integrates these signals directly:
+$A_{\text{total}}(s_{t},a_{t})=A_{1}+A_{2}(s_{t},a_{t}),$ (10)
+where $A_{1}$ represents outcome-based advantage and $A_{2}$ represents process-based advantage for token $s_{t}$ with action $a_{t}$.
+We then apply a binary mask to eliminate advantage values for non-code tokens:
+$\hat{A}(s_{t},a_{t})=A_{\text{total}}(s_{t},a_{t})\cdot\mathbbm{1}_{\text{is\_code}}(s_{t}).$ (11)
+This integrated advantage function balances sequence-level feedback with token-specific guidance.
+$A_{1}$ provides sequence-level guidance from watermark effectiveness and functional correctness, while $A_{2}$ delivers immediate, token-level feedback for watermark positioning and green token selection.
+Intuitively, our integrated reward system assigns high advantages to tokens that simultaneously maintain code functionality, contribute to effective watermark detection, and are selected from the green set.
+This approach guides our watermarked policy to identify optimal positions and select tokens that preserve functionality while introducing subtle, statistically detectable deviations.
+KL Regularization.
+To ensure that the watermarked policy $\pi_{\theta\oplus\phi}$ does not deviate excessively from natural code generation patterns, thereby maintaining code stealth and linguistic quality, we incorporate a Kullback-Leibler (KL) divergence penalty into the overall GRPO loss function.
+This term, $D_{\text{KL}}(\pi_{\theta\oplus\phi}\parallel\pi_{\text{ref}})$, regularizes the watermarked policy against the reference policy $\pi_{\text{ref}}$.
+Objective of GRPO.
+Combining all components discussed above, the final GRPO objective for our watermarked policy can be concisely expressed as:
+$\max_{\phi}$ $\mathbb{E}_{s\sim\mathcal{D}}\Bigg[\frac{1}{|s|}\sum_{t=1}^{|s|}\min\Bigg[\frac{\pi_{\theta\oplus\phi}(s_{t}|s_{<t})}{\pi_{\text{ref}}(s_{t}|s_{<t})}\hat{A}(s_{t},a_{t}),\hskip-1.99997pt\text{clip}\left(\frac{\pi_{\theta\oplus\phi}(s_{t}|s_{<t})}{\pi_{\text{ref}}(s_{t}|s_{<t})},1-\varepsilon,1+\varepsilon\right)\hat{A}(s_{t},a_{t})\Bigg]\Bigg]$
+$\hskip-1.99997pt-\beta D_{\text{KL}}(\pi_{\theta\oplus\phi}\parallel\pi_{\text{ref}}),$ (12)
+where $\mathcal{D}$ represents rollout data from watermarked policy, $\hat{A}(s_{t},a_{t})$ is our masked advantage function incorporating both outcome-based and process-based rewards, $\varepsilon$ is a clipping parameter that limits policy updates, and $\beta$ controls KL regularization.
+This objective balances watermark detectability, code functionality, and natural generation patterns while only optimizing watermark parameters $\phi$.
+5 Experiments
+Evaluation Overview.
+We evaluate CodeTracer across five critical dimensions: (i) code functionality, (ii) watermark detectability, (iii) robustness against attacks, (iv) efficiency and (v) transferability.
+Following Lee et al.
+(2023), we employ two primary metrics: Pass@k for functionality assessment and AUROC and TPR@5%FPR for detection performance measurement (Chen et al., 2021).
+Models and Datasets.
+We employ OpenCoder-1.5B-Instruct (Huang et al., 2024) as our primary backbone LLM and validate our findings on OpenCoder-8B-Instruct.
+We evaluate performance on three established benchmarks: HumanEval (Chen et al., 2021) and MBPP (Austin et al., 2021) for Python evaluation, and HumanEvalPack (Muennighoff et al., 2023) for cross-language evaluation.
+Baselines.
+We compare CodeTracer against two categories of baselines.
+For post-hoc detection methods, we evaluate against logp(x), LogRank (Gehrmann et al., 2019), DetectGPT (Mitchell et al., 2023), and GPTZero (Tian et al., 2023).
+For active watermarking approaches, we compare with WLLM (Kirchenbauer et al., 2023a), EXP-edit (Kuditipudi et al., 2023) and SWEET (Lee et al., 2023).
+Technical Implementation.
+Detailed implementation specifications, including dataset details, baseline configurations, model architectures, and training details are provided in Appendix C.
+We mainly follow the experiment settings of Lee et al.
+(2023) to adapt it to our setting.
+In accordance with established reinforcement learning practices, we initially apply supervised fine-tuning to $\pi_{\phi}$ using proxy signals to establish code distribution knowledge through next token prediction.
+The training process only needs to be performed once and can be completed in 1 day on a single A100 GPU.
+Table 1: Performance (%) comparison of detection methods. Bold: best performance for watermarks.
 Category Dataset HumanEval MBPP
-Method / Metric Pass@1 Pass@10 AUROC TPR Pass@1 Pass@10 AUROC TPR
+Method Pass@1 Pass@10 AUROC TPR Pass@1 Pass@10 AUROC TPR
 No Watermark Base 65.42 79.17 - - 43.35 51.65 - -
-Post-hoc Methods logp(x) 65.42 79.17 47.59 4.27 43.35 51.65 47.77 6.40
+Post-hoc logp(x) 65.42 79.17 47.59 4.27 43.35 51.65 47.77 6.40
 LogRank 65.42 79.17 47.66 1.82 43.35 51.65 48.76 7.80
@@
 GPTZero 65.42 79.17 52.00 5.50 43.35 51.65 41.10 2.80
-Oracle Watermarks SWEET 61.77 75.45 83.10 42.68 41.01 48.35 84.40 32.00
-Practical Watermarks WLLM 58.05 70.35 70.17 20.73 39.66 47.22 76.44 27.80
+Watermarks WLLM 58.05 70.35 70.17 20.73 39.66 47.22 76.44 27.80
 EXP-edit 59.29 72.41 66.50 25.61 40.16 50.25 51.22 10.60
-CodeTracer 60.82 79.72 82.95 46.34 41.45 51.15 79.69 22.60
-Integration into Advantage Computation.
-Our approach calculates advantages at two complementary levels to optimize the watermarked policy.
-First, we compute outcome advantages $A_{1}$ from the execution and watermark detection rewards ($R_{1}$ and $R_{2}$) using GRPO’s outcome reward methodology.
-Second, we derive process advantages $A_{2}$ from the token-level watermark reward ($R_{3}$), normalizing these values across all tokens in all rollouts for each prompt.
-Our framework categorizes tokens into four distinct types: green tokens (watermarked and included in set $G$), red tokens (watermarked but falling in set $R$), non-watermarked tokens (where decision $w=0$), and non-code text tokens.
-To maintain focus specifically on code watermarking, we exclude non-code tokens entirely from process advantage calculations.
-The final advantage function for each token $s_{t}$ integrates these signals directly:
-$A_{\text{total}}(s_{t},a_{t})=A_{1}+A_{2}(s_{t},a_{t}),$ (10)
-where $A_{1}$ represents outcome-based advantage and $A_{2}$ process-based advantage for token $s_{t}$ with action $a_{t}$.
-We then apply a binary mask to eliminate advantage values for any non-code tokens:
-$\hat{A}(s_{t},a_{t})=A_{\text{total}}(s_{t},a_{t})\cdot\mathbbm{1}_{\text{is\_code}}(s_{t}).$ (11)
-This integrated advantage function balances global sequence-level feedback with targeted token-specific guidance.
-While the outcome component $A_{1}$ evaluates tokens within the context of overall watermark effectiveness and functional code execution, the process component $A_{2}$ provides immediate direction for token selection decisions.
-Ideally, our integrated reward system assigns high advantage to tokens that simultaneously maintain code functionality, contribute to effective watermark detection, and are selected from the green set.
-This comprehensive approach guides our watermarked policy $\pi_{\theta\oplus\phi}$ to identify optimal watermarking positions and select tokens that preserve functionality while introducing subtle, statistically detectable deviations from normal token distributions.
-By selectively masking non-code tokens, we concentrate the learning signal exclusively on elements most relevant to code watermarking, substantially improving policy optimization efficiency.
-KL Regularization.
-To ensure that the watermarked policy $\pi_{\theta\oplus\phi}$ does not deviate excessively from natural code generation patterns, thereby maintaining code stealth and linguistic quality, we incorporate a Kullback-Leibler (KL) divergence penalty into the overall GRPO loss function.
-This term, $\beta\cdot D_{\text{KL}}(\pi_{\theta\oplus\phi}\parallel\pi_{\text{ref}})$, regularizes the watermarked policy against the reference policy $\pi_{\text{ref}}$.
-5 Experiments
-Evaluation Dimensions and Metrics.
-We evaluate CodeTracer across three critical dimensions: (i) code functionality, (ii) watermark detectability, and (iii) robustness against attacks.
-For code functionality, we employ the Pass@k metric [3].
-Watermark detection performance is measured by AUROC and TPR@5%FPR metrics.
-Comparative Baselines.
-For a comprehensive comparison, we compare CodeTracer against established post-hoc detection methods including logp(x), LogRank [5], DetectGPT [18], and GPTZero [21].
-Moreover, we compare with active watermarking approaches such as WLLM [9] and EXP-edit [11].
-We also include SWEET [13] serves as a reference baseline.
-However, to be noted, SWEET needs to redo the forward during the detection phase, which adopts a different setting from other baselines.
-A detailed discussion can be found in Appendix C.2.
-Technical Implementation.
-We primarily utilize OpenCoder-1.5B-Instruct [Huang2024OpenCoderTO] as the backbone LLM, with additional validation on the larger OpenCoder-8B-Instruct model.
-Performance is assessed on HumanEval [3] and MBPP [2] for Python code generation.
-Detailed implementation specifications, including dataset details, baseline configurations, model architectures, and training details are provided in Appendix C.
-In accordance with established reinforcement learning practices, we initially apply supervised fine-tuning to $\pi_{\phi}$ using proxy signals to establish fundamental code distribution knowledge through a combination of next token prediction and entropy distribution with binary cross entropy loss.
-Details and other techniques involved can be found in Appendix B.
-5.1 Results on Watermark Detection Performance and Code Functionality
-We present experiments evaluating CodeTracer’s performance on code functionality and watermark detection with HumanEval and MBPP in Table 1.
-CodeTracer outperforms other practical watermarking methods across most metrics.
-It shows minimal code quality degradation (Pass@1 scores of $60.82\%$ on HumanEval and $41.45\%$ on MBPP), while enabling effective detection (AUROC values of $82.95\%$ and $79.69\%$).
-With the highest TPR ($46.34\%$) on HumanEval among practical watermarks and detection performance comparable to oracle methods, CodeTracer emerges as the leading practical solution on these two datasets.
-5.2 Robustness Analysis
-Table 2: Robustness evaluation (%) under code modification attacks. Bold values indicate best performance.
+[…]
+[…]
+CodeTracer 62.65 77.11 77.71 32.32 42.10 52.01 78.42 31.60
+Figure 2: Training on 1.5B LLM and evaluating on 8B OpenCoder-8B-Instruct LLM.
+Table 2: Relative computational overhead of CodeTracer with a 1.5B base LLM. Our watermark model runs in parallel, so inference delay is only tensor slicing operation delay.
+Component Parameters Memory Delay
+[…]
+1.5B LLM $\pi_{\theta}$ 1,500M  3-6GB  500-800ms
+Relative Overhead < 10% < 15% < 0.02%
+5.1 Main Results on Benchmark Datasets.
+We present experiments evaluating CodeTracer’s performance on code functionality and watermark detection with HumanEval and MBPP in Table 1. We observe a clear improvement over other baselines.
+Post-hoc vs.
+Watermarks.
+Post-hoc detection methods, including logp(x), LogRank, DetectGPT, and GPTZero, preserve original generation but demonstrate consistently poor discrimination performance.
+Their AUROC values range from 47.59% to 52.00% on HumanEval and 41.10% to 48.76% on MBPP, close to random guessing (50%).
+This indicates that with the advancement of LLMs, post-hoc detection methods struggle to distinguish AI-generated code from human-written code.
+Code Functionality and Watermark Detectability.
+Watermarking methods demonstrate superior detection performance compared to post-hoc approaches, achieving AUROC values of 66.50-77.71% versus 47.59-52.00%.
+However, existing watermarking baselines suffer from noticeable code quality degradation, with Pass@1 scores dropping to 58.05-61.65% on HumanEval, representing a substantial trade-off between detection capability and functionality.
+CodeTracer achieves considerably better balance in this trade-off with the highest Pass@1 scores (62.65% on HumanEval, 42.10% on MBPP) and superior detection capability (AUROC of 77.71% and 78.42%), consistently outperforming SWEET and other baselines across both functionality and watermark detectability metrics.
+[…]
+00footnotetext: SWEET and SWEET use different entropy thresholds to explore detectability-functionality trade-offs.
+We adapt SWEET to use fixed context windows (vs.
+full generation history and original LLM) for practical and fair comparison with other baselines.
+See Appendix C.2 for details about entropy threshold setup.
+Table 3: Robustness evaluation (%) under code modification attacks. Bold values indicate best performance. Settings follow Lee et al. (2023).
 Attack Metric WLLM EXP-edit CodeTracer
@@
 TPR 20.12 9.76 29.11
-Figure 2: Performance for OpenCoder-8B-Instruct.
-Figure 3: The training dynamics for CodeTracer-0.
-As shown in Table 2, we evaluate CodeTracer’s robustness against DIPPER [10] paraphrasing and variable renaming attacks [13].
-While all methods experience performance degradation under attacks, CodeTracer consistently outperforms competing approaches.
-Against DIPPER attacks, CodeTracer maintains the highest detection capability (AUROC 58.42%, TPR 14.31%).
-Similarly, under variable renaming attacks, CodeTracer demonstrates superior resilience (AUROC 73.36%, TPR 29.11%), indicating its effectiveness in real-world scenarios with code modifications.
-5.3 Further Analysis
-Scalability to Larger Models.
-To verify the scalability or our approach, we evaluate CodeTracer on the larger OpenCoder-8B-Instruct model.
+Table 4: Ablation study on different reward components for CodeTracer training.
+Method Pass@1 (%) AUROC (%) TPR (%)
+CodeTracer 60.82 82.95 46.34
+w/o $A_{2}$ 61.15 (+0.33) 75.11 (-7.84) 30.29 (-16.05)
+w/o $A_{1}$ 60.34 (-0.48) 79.52 (-3.43) 34.91 (-11.43)
+Table 5: Performance of model variants.
+Method Pass@1 (%) AUROC (%) TPR (%)
+CodeTracer-1 62.65 77.71 32.32
+CodeTracer-2 60.82 82.95 46.34
+Figure 3: Cross-language evaluation on Java and C++. CodeTracer achieves consistent cross-language performance.
+Figure 4: Training dynamics for CodeTracer-1 with RL training.
+Transferability to Larger Models.
+To verify the scalability of our approach, we train the watermark model on a smaller 1.5B LLM and directly apply it as a plug-in module to the larger OpenCoder-8B-Instruct without retraining.
 Results are shown in Figure 2.
 CodeTracer achieves minimal code functionality degradation (71.77% Pass@1 vs.
-72.04% for Base) while significantly outperforming other practical watermarking methods in detection capability (78.69% AUROC compared to 65.90% for WLLM and 54.21% for EXP-edit).
-These results demonstrate CodeTracer’s effectiveness and generalizability across different parameter scales.
-Figure 4: Performance (%) comparison with variant
-Dataset Performance Metrics
-Method Pass@1 AUROC TPR
-CodeTracer 60.82 82.95 46.34
-CodeTracer-0 58.65 77.50 27.00
-Training from Scratch.
-To evaluate the effectiveness of pure reinforcement learning without prior knowledge, we introduce CodeTracer-0, a variant trained without SFT initialization.
-As shown in Figure 3, despite beginning with low detection rewards (approximately 0.3), the model consistently improves throughout RL training, eventually stabilizing around 0.8 for outcome-based watermark detection reward $R_{2}$, which is a proxy for z-score.
-Table 4 quantifies the final performance: CodeTracer-0 achieves reasonable results with Pass@1 of 58.65%, AUROC of 77.50%, and TPR of 27.00%, though not matching our fully-initialized CodeTracer.
-These results demonstrate that while good initialization enhances RL performance, our watermarking method remains effective even when trained entirely from scratch.
-Figure 5: Computational Efficiency Comparison
-Component Resource Metrics
-Method Parameters Memory Inference
-[…]
-Baseline 1.5B Model 1,500M $\sim$3-6GB $\sim$500-800ms
-Relative Overhead $10\% <$$15\% <$$0.02\%$
-Computation Efficiency
-[…]
+72.04% for Base) while significantly outperforming other watermarking methods in detection capability (78.69% AUROC compared to 65.90% for WLLM and 54.21% for EXP-edit).
+This demonstrates that our watermark model can serve as an effective plug-in module for larger models without requiring model-specific retraining.
+Robustness Analysis.
+Table 5.1 evaluates CodeTracer’s robustness against DIPPER (Krishna et al., 2024) paraphrasing and variable renaming attacks (Lee et al., 2023).
+While all methods degrade under attacks, CodeTracer consistently outperforms baselines.
+Against DIPPER attacks, CodeTracer maintains superior detection (AUROC 58.42%, TPR 14.31%), and under renaming attacks demonstrates resilience (AUROC 73.36%, TPR 29.11%), indicating effectiveness in real-world scenarios.
+Cross-Language Capabilities.
+Figure 4 shows CodeTracer’s cross-language performance on Java and C++.
+CodeTracer achieves strong detection while maintaining functionality comparable to baselines across languages, showing consistent effectiveness across different programming languages.
+5.2 Further Analysis
+Reward Components.
+Table 5.1 evaluates the impact of different reward components on CodeTracer’s performance.
+CodeTracer achieves optimal performance with both process and outcome rewards.
+Ablating either component leads to notable degradation, with removing the process reward causing larger detection drops while removing the outcome reward affects both performance.
+Performance Tradeoff.
+Table 5.1 demonstrates the controllable trade-off between code functionality and watermark detectability.
+CodeTracer-1 uses solely RL training without supervised fine-tuning initialization, achieving higher functionality (Pass@1 62.65%) but lower detection performance (AUROC 77.71%).
+Figure 4 shows the training dynamics of CodeTracer-1, demonstrating consistent improvement throughout RL training.
+In contrast, CodeTracer-2 incorporates both supervised fine-tuning and RL, trading some functionality for significantly improved detection capability.
 6 Conclusion
@@
 Our multi-component reward system balances statistical detectability with code integrity through execution feedback and token-level guidance.
-Evaluations demonstrate that CodeTracer significantly outperforms existing approaches in both detection accuracy and functionality preservation, offering a promising solution for identifying AI-generated code while maintaining its utility.
-7 Limitations
-While CodeTracer demonstrates significant advantages, certain limitations merit consideration.
-The reinforcement learning process for training the watermarking policy $\pi_{\theta}$, though optimized with rule-based rewards, can be computationally intensive, especially when applied to extensive codebases or a wide array of programming languages.
-The evolving nature of adversarial attacks also means that while CodeTracer enhances robustness, ongoing research will be crucial to maintain resilience against future sophisticated attempts at watermark evasion.
-The dynamic, per-token decision-making by the watermarked policy $\pi_{\theta\oplus\phi}$, including its determination of green/red list partitions, could also introduce a slight inference latency during code generation compared to simpler, non-adaptive watermarking methods; the impact of this would depend on the specific complexity of the learned policy.
-Finally, while our use of Gumbel Soft Top-k effectively enables differentiable learning for discrete vocabulary partitioning, the characteristics of this approximation and its potential interaction with highly diverse or unusual coding constructs represent an area for continued detailed analysis.
+Evaluations demonstrate that CodeTracer significantly outperforms existing approaches, offering a promising solution for identifying AI-generated code while maintaining the code functionality.
 Appendix A Watermark Generation and Detection Algorithm of CodeTracer
@@
 The algorithm modifies logits selectively based on these decisions, applying bias $\delta$ to green list tokens when $w=1$.
-Algorithm 1 CodeTracer Watermark Generation
 0: Prompt $\mathbf{x}$, Context window size $c$, Green list ratio $\gamma$, Bias strength $\delta$
@@
 3:  Compute base LLM logits $\mathbf{l}=\pi_{\theta}(\mathbf{ctx})$
-4:  Compute context window $\mathbf{ctx}=\text{concat}(\mathbf{x},\mathbf{\tilde{s}})[-(c):]$
+4:  Compute context window $\mathbf{ctx}=\text{concat}(\mathbf{x},\mathbf{\tilde{s}})[-(c)\mathrel{\mathop{\ordinarycolon}}$]
 5:  Compute watermark model outputs $(w_{\phi},\mathbf{l}_{\phi})=\pi_{\phi}(\mathbf{ctx})$
-6:  Compute watermark decision $w=\mathbbm{1}_{w_{\phi}>0.5}$
+6:  Compute watermark decision $w=\mathbbm{1}_{w_{\phi}>0}$
 7:  if $w=1$ then
@@
 11:   Set $\mathbf{\tilde{l}}=\mathbf{l}$ (no modification)
-12:  end if
 […]
-14:  Append $\tilde{s}_{t}$ to $\mathbf{\tilde{s}}$
-15:end for
-16:return  $\mathbf{\tilde{s}}$
+13:  Append $\tilde{s}_{t}$ to $\mathbf{\tilde{s}}$
+14:return $\mathbf{\tilde{s}}$
+Algorithm 1 CodeTracer Watermark Generation
 A.2 Watermark Detection
 Algorithm 2 details our statistical watermark detection procedure.
-Following the framework of [9], we employ hypothesis testing with a key modification, that testing is performed selectively only at positions likely to contain watermarks.
+Following the framework of Kirchenbauer et al.
+(2023a), we employ hypothesis testing with a key modification, where testing is performed selectively only at positions likely to contain watermarks.
 Given a code sequence $\mathbf{s}$, the detector first reconstructs the vocabulary partitions and switch probabilities using the same watermark model from generation.
-The detection employs a one-sided hypothesis test using the z-statistic to determine whether green tokens appear more frequently than random chance would predict, with a positive z-score above the threshold indicating watermark presence.
-Algorithm 2 CodeTracer Watermark Detection
+Detection employs a one-sided hypothesis test using the z-statistic to determine whether green tokens appear more frequently than random chance would predict, with a positive z-score above the threshold indicating watermark presence.
 0: Code sequence $\mathbf{s}$, Context window size $c$, Green list ratio $\gamma$, Detection threshold $\tau$
@@
 3:for each position $t$ in $\mathbf{s}$ do
-4:  Compute context window $\mathbf{ctx}=\mathbf{s}[\max(0,t-c):t]$
+14:  Compute context window
+$\mathbf{ctx}=\mathbf{s}[\max(0,t-c)\mathrel{\mathop{\ordinarycolon}}t]$
 5:  Compute watermark model outputs $(w_{\phi},\mathbf{l}_{\phi})=\pi_{\phi}(\mathbf{ctx})$
@@
 11:    $\text{green\_tokens}++$
-12:   end if
-13:   Add $t$ to positions_checked
-14:  end if
-15:end for
-16: Compute z-score: $z=\frac{\text{green\_tokens}-\text{watermarked\_positions}\cdot\gamma}{\sqrt{\text{watermarked\_positions}\cdot\gamma\cdot(1-\gamma)}}$
-17:if $z>\tau$ then
-18:  return  “watermarked”
-19:else
-20:  return  “not watermarked”
-21:end if
+12:   Add $t$ to positions_checked
+213: Compute z-score:
+$z=\frac{\text{green\_tokens}-\text{watermarked\_positions}\cdot\gamma}{\sqrt{\text{watermarked\_positions}\cdot\gamma\cdot(1-\gamma)}}$
+14:if $z>\tau$ then
+15:  return “watermarked”
+16:else
+17:  return “not watermarked”
+Algorithm 2 CodeTracer Watermark Detection
 Appendix B Advanced Training And Generation Techniques
@@
 While our reinforcement learning (RL) algorithm can train from random initialization, we find that supervised fine-tuning (SFT) provides a significantly more effective starting point.
-Since watermarked code is unavailable for direct supervision, we instead use entropy and next token prediction of clean code as a proxy.
-During SFT, the watermark model $\pi_{\phi}$ is trained on two complementary objectives: predicting watermark decisions $w_{\phi}$ to learn entropy distribution patterns across code structures, and optimizing logits $l_{\phi}$ to capture the underlying distribution of tokens in code sequences.
+Since watermarked code is unavailable for direct supervision, we instead use entropy and next-token prediction of clean code as a proxy.
+During SFT, the watermark model $\pi_{\phi}$ is trained on two complementary objectives: predicting watermark decisions $w_{\phi}$ to learn entropy distribution patterns across code structures, and optimizing logits $l_{\phi}$ to capture the underlying token distribution in code sequences.
 This dual-objective approach establishes a strong foundation that encodes both code syntax patterns and appropriate watermarking signals before RL optimization begins.
@@
 For less certain positions, the model learns to disable watermarking ($w=0$), which receives a neutral reward.
-This approach significantly reduces false positive detection rates while maintaining high true positive rates.
+This approach significantly reduces false-positive detection rates while maintaining high true-positive rates.
+Table 6: Comprehensive hyperparameter settings for the CodeTracer watermarking framework.
+The parameters are organized by their functional categories: watermark configuration controlling detectability, model architecture defining the watermark policy network structure, training configuration governing the optimization process, and reward configuration balancing the various objectives.
 Parameter Value Description
@@
 Temperature 1.0 Sampling temperature for generation
-Table 3: Comprehensive hyperparameter settings for the CodeTracer watermarking framework.
-The parameters are organized by their functional categories: watermark configuration controlling detectability, model architecture defining the watermark policy network structure, training configuration governing the optimization process, and reward configuration balancing the various objectives.
 B.5 Limitations of Using LLMs as Watermark Policy Networks
@@
 Table 6 presents the comprehensive hyperparameter configuration for our CodeTracer framework, categorized by functional components to facilitate reproducibility and highlight critical design decisions across our experimental setup.
-For paper writing, we use AI tools to refine the overall expression.
 C.1 Datasets
 We evaluate on standard code generation benchmarks that span multiple programming languages and tasks.
-The primary evaluation uses HumanEval [3] and MBPP [2], which provide comprehensive Python programming challenges with associated test cases and reference implementations.
+The primary evaluation uses HumanEval (Chen et al., 2021) and MBPP (Austin et al., 2021), which provide comprehensive Python programming challenges with associated test cases and reference implementations.
 C.2 Baselines
 We compare against two categories of methods: post-hoc detection and active watermarking.
-Post-hoc detection methods preserve original generation and include zero-shot approaches: logp(x), LogRank [5], DetectGPT [18], and GPTZero [21].
-Active watermarking methods include WLLM [9] and EXP-edit [11], which operate under the practical constraints.
-We also include SWEET [13] as a reference, though comparisons are not fair since it requires access to the original LLM and prompts for entropy calculation.
+Post-hoc detection methods preserve original generation and include zero-shot approaches: logp(x), LogRank (Gehrmann et al., 2019), DetectGPT (Mitchell et al., 2023), and GPTZero (Tian et al., 2023).
+Active watermarking methods include WLLM (Kirchenbauer et al., 2023a) and EXP-edit (Kuditipudi et al., 2023), which operate under the practical constraints.
+We also include SWEET (Lee et al., 2023), though comparisons are not fair since it requires access to the original LLM and prompts for entropy calculation.
+We adapt SWEET to use a practical setting for fair comparison with other baselines.
 We thoroughly tune the hyperparameters for each baseline following their setup.
@@
 For the DetectGPT implementation, we used T5-3B as our model.
-Following the original DetectGPT paper and SWEET [18, 13] , we set the span length to 2 words and applied masking to 20% of the text.
+Following the original DetectGPT paper and SWEET (Mitchell et al., 2023; Lee et al., 2023) , we set the span length to 2 words and applied masking to 20% of the text.
 For each test, we generated 100 perturbations to ensure robust detection.
@@
 For the SWEET baseline implementation, we conducted a systematic hyperparameter search for the entropy threshold, exploring values from 0.3 to 1.2 with increments of 0.3.
-Our experiments revealed optimal performance with an entropy threshold of 1.2 for the HumanEval dataset and 0.3 for MBPP.
-In detection staga, SWEET needs to use original LLM, prompt and complete generation sequence to compute the entropy.
+Our experiments revealed optimal performance with an entropy threshold of 1.2 and 0.9 for the HumanEval dataset and 0.3 and 0.6 for MBPP.
+In detection stage, SWEET needs to use original LLM, prompt and complete generation sequence to compute the entropy.
 This requirement severely restricts their practical applicability in real-world scenarios.
 For practical deployment, an ideal solution would depend solely on the code snippet itself, as access to the generation context is typically unavailable.
-Thus, we only use SWEET as a reference model in our performance comparision.
+Thus, we adapt SWEET to use fixed context windows (vs.
+full generation history and original LLM) for practical and fair comparison with other baselines.
+We select two entropy thresholds for SWEET to demonstrate the trade-off between watermark detectability and code functionality.
 EXP-edit.
-We evaluate EXP-edit [11] following the methodology of [13].
+We evaluate EXP-edit (Kuditipudi et al., 2023) following the methodology of (Lee et al., 2023).
 In our experiments, we set temperature=$0.2$ and top-p=$0.95$.
@@
 We implement reinforcement learning training using GRPO with several customizations for the code watermarking task.
-Our training pipeline processes the mix of open-r1/verifiable-coding-problems-python dataset, HumanEval dataset and MBPP dataset.
+Our training pipeline processes the mix of prompts from open-r1/verifiable-coding-problems-python dataset, HumanEval dataset and MBPP dataset.
 The motivation is to make sure model be familiar the distribution of these dataset instructions, which are representative code examples for code generation tasks.
-The watermark policy model optimization uses a learning rate of $1.0\times 10^{-5}$ with a cosine scheduler including a minimum learning rate rate of 0.1.
+The watermark policy model optimization uses a learning rate of $1.0\times 10^{-5}$ with a cosine scheduler including a minimum learning rate of 0.1.
 We implement a short warmup period (3% of total steps) to stabilize early training.
@@
 This multi-faceted security approach provides robust protection against both detection evasion and watermark forgery in practical deployment scenarios.
+Appendix E Use of Large Language Models
+Large Language Models (LLMs) were used solely as a general-purpose writing assistant to check grammar and improve expression clarity throughout this paper.
+The LLMs did not contribute to research ideation, methodology development, experimental design, or generation of novel content.
+All technical contributions, insights, and findings presented in this work are entirely the result of the authors’ original research.